### v1

# Rethinking Softmax: Self-Attention with Polynomial Activations

Abstract
This paper challenges the conventional belief that softmax attention in transformers is effective primarily because it generates a probability distribution for attention allocation. Instead, we theoretically show that its success lies in its ability to implicitly regularize the Frobenius norm of the attention matrix during training. We then explore alternative activations that regularize the Frobenius norm of the attention matrix, demonstrating that certain polynomial activations can achieve this effect, making them suitable for attention-based architectures. Empirical results indicate these activations perform comparably or better than softmax across various computer vision and language tasks, suggesting new possibilities for attention mechanisms beyond softmax.
1 Introduction
Transformer architectures (Vaswani et al., 2017) have become the state-of-the-art model architecture for diverse areas such as natural language processing (NLP) (Vaswani et al., 2017; Devlin et al., 2018; Zhuang et al., 2021; Zhen et al., 2022), computer vision (Dosovitskiy et al., 2020; Carion et al., 2020; Liu et al., 2021; Touvron et al., 2021), and robotics (Fu et al., 2024; Maiti et al., 2023; Salzmann et al., 2020). A key component in the transformer architecture is the softmax attention block, enabling transformers to evaluate the importance of individual input elements during output generation. This feature facilitates an efficient method to attend to diverse input elements throughout training, allowing transformers to effectively capture spatial dependencies within sequential data. Unlike traditional recurrent neural networks (RNNs) and convolutional neural networks (CNNs), transformers exhibit the ability to scale to large datasets without a significant degradation in performance. This characteristic has made them an ideal architecture for handling large-scale machine learning tasks.
Softmax is widely recognized for its effectiveness in attention mechanisms due to its ability to produce an attention matrix that meets three key conditions (Vaswani et al., 2017; Dosovitskiy et al., 2020; Zhen et al., 2022; AUEB et al., 2016): (i) non-negativity, (ii) rows that are normalized to sum to 1 and (iii) sparsity. The general consensus is that non-negativity guarantees that attention weights remain positive, facilitating the model in assigning significance to various input elements. The normalization constraint ensures that the attention weights for all input elements collectively sum to $1$, rendering the weights interpretable as probabilities. Additionally, sparsity aids the model in focusing on a select few key elements, thereby enhancing efficiency. It has been argued that these properties are crucial for enabling the attention mechanism to attend to pertinent segments of the input sequence while efficiently filtering out irrelevant details. However, this approach to attention has become somewhat axiomatic as it is mostly motivated by empirical results with little theoretical foundation. Despite the exploration of alternative activations in several studies (Shen et al., 2023; Fang et al., 2022; Correia et al., 2019), softmax attention continues to dominate, largely due to its interpretability.
In this paper, we question this view by proposing that the effectiveness of softmax stems from its implicit regularization of the Frobenius norm of the attention matrix during training, preventing attention weights from becoming excessively large or small. We then derive a theoretical framework that produces polynomial activations that deliberately violate one or more of the three conditions mentioned earlier, yet are able to regularize the Frobenius norm of the attention weights during training. Our findings demonstrate that such activations can achieve comparable or even superior performance to softmax across various vision and natural language processing (NLP) tasks, even though they seem to violate our understanding of attention.
We advise the reader that this paper diverges from the usual pursuit of creating cutting-edge transformer architectures for achieving state-of-the-art results on benchmark datasets. Instead, our focus is on critically examining softmax attention to determine whether its effectiveness is a result of true interpretability or a more nuanced regularization mechanism. By questioning established views, we aim to uncover deeper insights into transformer architectures that could lead to broader applications and improved understanding. Nonetheless, we validate our theory on multiple transformer-based tasks, including image classification, segmentation, object detection, and NLP, often achieving results that match or surpass those of softmax attention. Our main contributions are:
  • 1.
    We question the widely accepted notion that softmax’s effectiveness in attention mechanisms is solely due to its ability to produce normalized sparse attention weights. Instead, we theoretically show that softmax has a regularization effect on attention and argue this plays a crucial role in its success.
  • 2.
    We explore activations that deliberately deviate from traditional softmax attention conditions. These activations are found to regularize the Frobenius norm of the attention matrix during training, akin to softmax, and demonstrate comparable or superior performance across various vision and NLP tasks.
2 Related Work
Several studies have explored alternative activations for attention mechanisms in transformers. Shen et al. (2023) investigated ReLU activations, finding them to outperform softmax in tasks with long sequences, such as document translation. Banerjee et al. (2020) examined Taylor series approximations to softmax, which showed superior performance to softmax in image classification. Wang et al. (2021) proposed periodic alternatives to softmax, designed to provide better gradients for attention mechanisms and achieved better results than softmax on simple networks for image classification. Koohpayegani & Pirsiavash (2024) demonstrated that applying $l^{1}$ normalization to linear attention mechanisms can achieve performance comparable to softmax. Our work differs from all these in that we identify a clear theoretical relationship between the scale of the Frobenius norm of the self-attention matrix and the input sequence length. Using this insight to derive potential acitvations that can perform on par with softmax.
3 Preliminaries and Notation
In this section we outline the definition of a transformer via the transformer block and set the notation of various mathematical quantities we will be using in future sections. For more details on transformers the reader can consult Vaswani et al. (2017); Dosovitskiy et al. (2020).
Transformer architectures comprise of transformer blocks, defined as follows. A transformer block is a mapping $\mathbf{T}:\mathbb{R}^{N\times D}\rightarrow\mathbb{R}^{N\times D}$ defined as
$\mathbf{T}(x)=\mathbf{F}(\mathbf{A}(x)+x)$ (3.1)
where $\mathbf{F}$ is a feedforward MLP with a residual connection and $\mathbf{A}$ is an attention head.
The attention head $\mathbf{A}$ is defined as follows: It comprises of three learnable matrices, a query ($q$), key ($k$) and value ($v$) defined by: $q=QX$, $k=KX$, $v=VX$ for an input sequence $X\in\mathbb{R}^{N\times D}$ with $Q$, $K\in\mathbb{R}^{D\times d}$ and $V\in\mathbb{R}^{D\times M}$. The attention head $\mathbf{A}(X)$ is then defined by
$\mathbf{A}(X)=\phi(\mathcal{S}(q,k))v$ (3.2)
where $\mathcal{S}$ is a similarity transformation and $\phi$ is an activation function. The most common used $\mathcal{S}$ is the dot-product: $\mathcal{S}(q,v)=qk^{T}$, known as self-attention, and will be the one we focus on in this paper. The most common activation function $\phi$ that is used by authors is softmax. This leads to the most common form of the attention head given by
$\mathbf{A}(X)=\mathbf{softmax}\bigg{(}\frac{qk^{T}}{\sqrt{d}}\bigg{)}v=\mathbf{softmax}\bigg{(}\frac{XQK^{T}X^{T}}{\sqrt{d}}\bigg{)}XV.$ (3.3)
The function $\mathbf{softmax}$ is the matrix softmax map that applies the usual softmax function row-wise:
$\mathbf{softmax}\bigg{(}\begin{bmatrix}x_{11}&\cdots&x_{1n}\\ \vdots&\vdots&\vdots\\ x_{n1}&\cdots&x_{nn}\end{bmatrix}\bigg{)}=\begin{bmatrix}\frac{e^{x_{11}}}{\sum_{j=1}^{n}e^{x_{1j}}}&\cdots&\frac{e^{x_{1n}}}{\sum_{j=1}^{n}e^{x_{1j}}}\\ \vdots&\vdots&\vdots\\ \frac{e^{x_{n1}}}{\sum_{j=1}^{n}e^{x_{nj}}}&\cdots&\frac{e^{x_{nn}}}{\sum_{j=1}^{n}e^{x_{nj}}}\end{bmatrix}$ (3.4)
The factor $\frac{1}{\sqrt{d}}$, as explained in Vaswani et al. (2017), is a scaling to prevent the gradients of softmax from being too small. For the theoretical analysis in this paper we will only use the dot-product similarity $qk^{T}$ and call the $N\times N$ matrix $softmax(qk^{T})$ the softmax self-attention matrix. In the experiments section, Sec. 5, we will empirically validate our theoretical framework on more general softmax attention blocks.
For general transformer architectures, multiple heads $\mathbf{A}_{i}$ for $1\leq i\leq n$ are used. Each attention head is defined by equation 3.3 and then all outputs of each attention head are concatenated together before going into the feedforward layer.
We will need notation for the derivative of the matrix softmax map defined by equation 3.4. Given a matrix $A\in\mathbb{R}^{N\times N}$ we can differentiate the matrix map $\mathbf{softmax}$ at $A$ and obtain the gradient linear map $\mathbf{\nabla softmax}(A):\mathbb{R}^{N\times N}\rightarrow\mathbb{R}^{N\times N}$ that is defined by the formula
$\mathbf{\nabla softmax}(A):=\mathbf{Jsoftmax}(A)^{T}$ (3.5)
where $\mathbf{Jsoftmax}(A)$ is the Jacobian of $\mathbf{softmax}$ at $A$.
Given a matrix $A\in\mathbb{R}^{n\times m}$, we denote its Frobenius norm by $||A||_{F}$. Additionally, we use the notation $\mathbb{E}$ to represent the expectation of a random variable, where the specific random variable being considered will be clear from the context.
4 Theoretical Analysis
4.1 Implicit regulatization of Softmax
This section presents a theoretical result showing that the softmax activation imposes control over the Frobenius norm of the self-attention matrix in a way that grows sub-linearly with the input sequence’s token length. Additionally, we demonstrate that the gradient of the softmax with respect to the self-attention matrix also exhibits a similar degree of regularity. While previous work has analyzed the regularity of softmax self-attention through the lens of the Lipschitz constant (Kim et al., 2021; Castin et al., 2023), our theorem offers a novel perspective by directly linking the Frobenius norm regularity to the token length. This provides insights into how self-attention activations should scale with token length to maintain stability during training, especially with gradient descent-based algorithms.
Theorem 4.1.
Let $\mathbf{softmax}:\mathbb{R}^{N\times N}\rightarrow\mathbb{R}^{N\times N}$ be the matrix softmax map defined by equation 3.4 and let $\mathbf{\nabla softmax}(A):\mathbb{R}^{N\times N}\rightarrow\mathbb{R}^{N\times N}$ denote the gradient of $\mathbf{softmax}$ at $A\in\mathbb{R}^{N\times N}$. We then have the following bounds on the Frobenius norms
$||\mathbf{softmax}(A)||_{F}$ $\leq\sqrt{N}$ (4.1)
$||\mathbf{\nabla softmax}(A)||_{F}$ $\leq 2\sqrt{N}.$ (4.2)
The key implication of theorem 4.1 is that during the training of a transformer with softmax self-attention, the Frobenius norm of each softmax self-attention matrix remains bounded by a value that grows as $\mathcal{O}(\sqrt{N})$. This ensures that backpropagation through the weights of the self-attention matrix does not lead to excessively large gradients. The proof hinges on the fact that the row normalization inherent in softmax effectively controls the Frobenius norm. For a detailed proof see appendix A.1.1.
4.2 Polynomial activations for self-attention
In section 4.1, we demonstrated that softmax implicitly regularizes the Frobenius norm of the self-attention matrix. Building on this, we now show that by scaling specific polynomial activations, a similar regularization effect on the Frobenius norm can be achieved in expectation, closely replicating the impact of softmax.
Theorem 4.2.
Let $X\in\mathbb{R}^{N\times D}$ and $Q$, $K\in\mathbb{R}^{D\times d}$ be i.i.d random variables distributed according to $X\sim\mathcal{N}(0,\sigma_{x})$ and $Q$, $K\sim\mathcal{N}(0,\sigma_{t})$. We have the following expectations of the Frobenius norms of powers of the $N\times N$ matrix $(XQK^{T}X^{T})^{p}$ for $p\geq 1$
$\mathbb{E}\bigg{|}\bigg{|}\bigg{(}\frac{XQK^{T}X^{T}}{\sqrt{d}}\bigg{)}^{p}\bigg{|}\bigg{|}_{F}\leq\mathcal{O}(N)$ (4.3)
By scaling such an activation by $\frac{1}{\sqrt{N}}$ we can obtain a $\mathcal{O}(\sqrt{N})$ bound.
Corollary 4.3.
Assume the same conditions as in theorem 4.2. Then
$\mathbb{E}\bigg{|}\bigg{|}\frac{1}{\sqrt{N}}\bigg{(}\frac{XQK^{T}X^{T}}{\sqrt{d}}\bigg{)}^{p}\bigg{|}\bigg{|}_{F}\leq\mathcal{O}(\sqrt{N}).$ (4.4)
Corollary 4.3 establishes that activations of the form $\phi(x):=\frac{1}{\sqrt{N}}x^{p}$ provide a level of regularization, in expectation, similar to that of softmax when applied to the self-attention matrix. The proof of theorem 4.2 can be found in appendix A.1.2. The next property we want to prove is one similar to the gradient bound obtained in theorem 4.1. Since the self-attention matrix has parameters given by the queries $Q$ and keys $K$ (Vaswani et al., 2017), this implies that during the training of a transformer the $Q$ and $K$ matrices are the only aspects of the self-attention matrix that get updated. Therefore, we compute a regularity result with respect to the $Q$ and $K$ derivatives.
Theorem 4.4.
Let $X\in\mathbb{R}^{N\times D}$ and $Q$, $K\in\mathbb{R}^{D\times d}$ be i.i.d random variables distributed according to $X\sim\mathcal{N}(0,\sigma_{x})$ and $Q$, $K\sim\mathcal{N}(0,\sigma_{t})$. Then the expectation of the of the derivative of the matrix $\frac{(XQK^{T}X^{T})^{p}}{\sqrt{d}}$ w.r.t the $Q$ parameter matrix for $p\geq 1$ is given by
$\mathbb{E}\bigg{|}\bigg{|}\frac{\partial}{\partial Q}\bigg{(}\frac{(XQK^{T}X^{T})^{p}}{\sqrt{d}}\bigg{)}\bigg{|}\bigg{|}\leq\mathcal{O}(N)$ (4.5)
The above theorem then suggests that if we scale the polynomial $x\rightarrow x^{p}$ by $\frac{1}{\sqrt{N}}$ the $Q$ derivative will have $\mathcal{O}(\sqrt{N})$ growth.
Corollary 4.5.
Assume the same condition as in theorem 4.4. Then
$\mathbb{E}\bigg{|}\bigg{|}\frac{1}{\sqrt{N}}\frac{\partial}{\partial Q}\bigg{(}\frac{(XQK^{T}X^{T})^{p}}{\sqrt{d}}\bigg{)}\bigg{|}\bigg{|}\leq\mathcal{O}(\sqrt{N}).$ (4.6)
An analogous estimate holds for derivatives with respect to the $K$ matrix. The proof of theorem 4.4 can be found in appendix A.1.2.
Corollaries 4.3 and 4.5 suggest that polynomial activations of the form $\phi(x)=\frac{1}{\sqrt{N}}x^{p}$, with $p>0$, can achieve performance comparable to softmax when applied to self-attention matrices. In section 5, we empirically compare these activations to softmax and observe that they outperform softmax on a variety of transformer tasks. We focus on $p=1$ and $p=3$ as these polynomials clearly violate key aspects of softmax based attention, such as normalized rows, positivity, and sparsity. For larger values of $p$, performance declines due to the functions $\phi(x)=\frac{1}{\sqrt{N}}x^{p}$ having smaller gradients around 00 when $p$ is large, causing difficulties in training.
5 Experiments
In this section, we validate the theory from section 4 on a variety of transformer tasks. We perform the empirical validation on two primary activations from section 4, namely a cubic polynomial activation $\phi(x)=x^{3}$ and a linear polynomial $\phi(x)=x$. The goal will be to show that by suitably scaling these activations using the theory in section 4, we can achieve competitive performance when compared to softmax. For the rest of this section we will simply denote these activations by $x^{3}$ and $x$.
Figure 1: Training ViT-Tiny with the activation $\phi(x)=x^{3}$ with different sequence lengths and different scales. As the sequence length gets smaller, the log scale needed to obtain good accuracy decreases validating the theory from section 4.2.
5.1 Image classification
5.1.1 ViT-Tiny on Tiny-Imagenet:
In this section we test the theory from section 4 on the ViT-Tiny architecture (Steiner et al., 2021) trained from scratch on the Tiny-Imagenet dataset (Le & Yang, 2015).
Our first experiment was to test how the Top-$1\%$ accuracy changes for a ViT-Tiny trained on Tiny-Imagenet as we change the sequence length of the input and the scale predicted in corollaries 4.3 and 4.5 when using the activation $x^{3}$. According to the theory developed in section 4, the Frobenius norm scales according to $\mathcal{O}(\sqrt{N})$ when we scale $X^{3}$ by $\frac{1}{\sqrt{N}}$. Thus as the sequence length decreases we should see the amount of scaling in a Log scale decrease.
Figure 1 shows the results of this experiment. We considered four different input sequence lengths of sizes $256$, $64$, $16$ and $8$. We ran several ViT-Tiny architectures with a variety of scalings of the form $\mathcal{O}(\frac{1}{\sqrt{N}})$ where $N$ ranged below to above the sequence length. As can be seen from figure 1 as the sequence length got smaller the amount of scaling, shown in Log scale on the x-axis, needed for good accuracy got smaller verifying the theory in section 4.2.
The second experiment compared activations $x^{3}$ and $x$, along with scaled versions $\frac{1}{16}x^{3}$ and $\frac{1}{16}x$, against softmax using the Tiny-ViT architecture on Tiny-Imagenet. With a sequence length of $256$ ($\sqrt{256}=16$), we decided to take $\frac{1}{16}$ as the scale of the polynomial activations. The experiment used a patch size of 4, 3 attention heads, and 12 layers as described in Steiner et al. (2021). Results in table 1 show $\frac{1}{8}x^{3}$ outperforming softmax, while the unscaled version performed poorly. Similarly, $\frac{1}{16}x$ performed competitively with a significant drop in performance without scaling.
Figure 2 displays the Frobenius norm of the self-attention matrix during training for five activations in layers 2 and 12 of ViT-Tiny, averaged across all heads. Norms for $x^{3}$ and $x$ are higher than softmax, but scaling by $\frac{1}{16}$ reduces them to more stable levels, improving training stability. Similarly, figure 3 shows the Jacobian’s Frobenius norm, where scaling also brings the norms closer to softmax, ensuring more stable gradients. Further plots for other layers are in appendix A.2.2.
softmax $\frac{x^{3}}{16}$ $x^{3}$ $\frac{x}{16}$ $x$
Top-1% accuracy 50.26 50.5 45.3 47.9 31.78
Table 1: Comparison of Top-1% accuracy on Tiny-Imagenet between softmax and polynomial activations. The cubic activation outperforms softmax when the right scale of $\frac{1}{8}$ is applied. Similarly, the linear activation is competitive only with an optimal scale.
Figure 2: Frobenius norm of the self-attention matrix with five different activations in layer 2 and 12 of the ViT-Tiny architecture during training.
Figure 3: Frobenius norm of the Jacobian of the self-attention matrix with five different activations in layer 2 and 12 of the ViT-Tiny architecture during training.
5.1.2 Larger vision transformers on ImageNet-1k
For this experiment we carried out an image classification task using a variety of different vision transformers from the literature on the ImageNet-1k dataset. We found that the scales $\frac{1}{14}$ worked best for both $x^{3}$ and $x$.
We train all models on the ImageNet-1k dataset from scratch and report Top-1 accuracy on the validation set. We use PyTorch Paszke et al. (2019) and Timm Wightman (2019) libraries to train our models with similar setups to He et al. (2022) Liu et al. (2021). We examined our approach along with the following three transformer architectures to show its generalization:
- ViT: Dosovitskiy et al. (2020) is the pioneering work that interprets an image as a sequence of patches and processes it by a standard Transformer encoder as used in NLP. This simple, yet scalable, strategy works surprisingly well when coupled with pre-training on large datasets. We use ViT-Small which has the following settings: patch size = 16, embedding dimensions = 384, number of heads = 6, and layers = 12. Also, We use ViT-Base which has the following settings: patch size = 16, embedding dimensions = 768, number of heads = 12, and layers = 12.
- DeiT: Touvron et al. (2021) is a well-known transformer based on ViT. It is very similar to ViT except that it converged faster due to better training recipe since we do not use the distillation token imposed in DeiT. We use DeiT-Small which has the following settings: patch size = 16, embedding dimensions = 384, number of heads = 6, and layers = 12. Also, We use Deit-Base which has the following settings: patch size = 16, embedding dimensions = 768, number of heads = 12, and layers = 12.
- Swin Transformer: Liu et al. (2021) produces a hierarchical feature representation and proposes the shifted window based self-attention which is shown to be effective and efficient on vision problems. We use Swin-Small with 96 channels and Swin-Base with 128 channels in the hidden layers of the first stage. The window size is set to $M$ = 7 by default, the query dimension of each head is $d$ = 32, and layer numbers are ${2,2,18,2}$ for all experiments.
- XciT: Xiong et al. (2021) is a vision transformer architecture consisting of two different components compared to the standard ViT. Firstly, it has Local Patch Interaction in each block, which includes one depth-wise 3×3 convolution followed by Batch Normalization, GELU, and another depth-wise 3×3 convolution. Secondly, it uses Cross-Covariance attention, where the attention map is derived from the cross-covariance matrix computed over the key and query projections of the token features. We use XCiT-S12 with a patch size of 16 and XCiT-M24 with a patch size of 24.
The results are shown in table 2. The activation $\frac{1}{14}x^{3}$ performed best on the ViT’s and the Swin transformers while softmax performed best on the DeiT architectures. Further ablations with different scales and activations can be fond in appendix A.2.1.
Activation Models
ViT-Base ViT-Small DeiT-Base DeiT-Small Swin-Base Swin-Small XciT-Medium XciT-Small
softmax 79.6 80.2 78.9 79.6 83.0 83.3 81.1 81.2
$\frac{x^{3}}{14}$ 79.6 80.5 77.4 78.3 83.2 83.4 81.2 82.1
$x^{3}$ 77.8 78.6 76.1 76.3 79.7 79.9 78.1 78.3
$\frac{x}{14}$ 76.9 77.8 79.6 79.8 79.4 79.5 79.2 79.3
$x$ 73.2 73.9 77.7 77.9 77.8 77.9 76.4 76.6
Table 2: Comparsions of pre-training models with different activation functions on ImageNet-1k datasets. We report the classification top-1 accuracy (%).
Figure 4 plots the Frobenius norm of the self-attention matrix with the five different activations during training in layers 2 and 12, averaged over all heads within each layer, of the ViT-Small architecture. By scaling the activations $x^{3}$ and $x$ by $\frac{1}{14}$ we were able to control the scale of the Frobenius norm of the self-attention matrix and obtain scales comparable to softmax’s scale. Simiarly, figure 5 plots the Frobenius norm of the Jacobian of the self-attention matrix during training for layers 2 and 12, averaged over all heads. By scaling the activations $x^{3}$ and $x$ by $\frac{1}{14}$ we were able to control the scale of the Jacobian norm and obtain scales comparable to softmax’s scale during training. Plots for other layers of the architecture during training are given in appendix A.2.2.
Figure 4: Frobenius norm of the self-attention matrix with five different activations in layer 2 and 12 of the ViT-Tiny architecture during training.
Figure 5: Frobenius norm of the Jacobian of the self-attention matrix with five different activations in layer 2 and 12 of the ViT-Tiny architecture during training.
5.1.3 Visualizing self-attention with ViT-Base
We plotted heat maps of self-attention matrices using the $\frac{1}{14}x^{3}$ and softmax activations across two layers and heads after convergence, averaging over a fixed training batch of size 128. Figure 6 shows layer 2, head 8, highlighting the differences in attention patterns, with $\frac{1}{14}x^{3}$ containing both positive and negative values. Similarly, Figure 7 for layer 12, head 6 shows distinct patterns for each activation. Overall, the ViT-Base architecture with $\frac{1}{14}x^{3}$ exhibits notably different self-attention patterns compared to softmax.
We visualized how the self-attention matrix targets different regions of an image. Using an image from the ImageNet-1k validation set, we extracted the class token, reshaped it into a $(14,14)$ grid representing 196 patches, and then mapped it back to the original image size with nearest neighbor interpolation. Figure 8 shows the input image, while figure 9 illustrates the self-attention matrices for different activations in layer 12, head 6 of the ViT-Base architecture after convergence, highlighting their distinct focus areas.
Figure 6: Heat maps of the self-attention matrix in layer 2, head 8, of a ViT base architecture, comparing $\frac{1}{14}x^{3}$ (left) and softmax (right) activations after training. The stark difference in self-attention patterns between the two activations is evident, showing distinct distributions across input tokens.
Figure 7: Heat maps of the self-attention matrix in layer 12, head 6, of a ViT base architecture, comparing $\frac{1}{14}x^{3}$ (left) and softmax (right) activations after training. The contrast in self-attention patterns between the two activations is clearly visible.
Figure 8: Fish image from validation set of ImageNet-1k broken up into $196=14\times 14$ patches.
Figure 9: Comparing how $\frac{1}{14}x^{3}$ and softmax self-attention matrices focus on different parts of the fish image from figure 8 in layer 12, head 6, after the model has converged.
5.2 Object Detection and Instance Segmentation
In this section, in order to examine the transfer learning ability of our models, we demonstrate our approach to object detection and segmentation tasks by fine-tuning our ImageNet-pretrained XCiT model on them. Our experiments are conducted on COCO 2017Lin et al. (2014), which has 118K training images and 5K validation images with 80 categories. We integrate the XCiT architecture as the backbone in the Mask R-CNN (He et al., 2017) detector with a Feature Pyramid Network (FPN). Due to XCiT’s inherently columnar design, we adapt it for FPN compatibility by extracting features from various layers for XCiT-S12. These features have a consistent stride of either 16. The feature resolutions are then adjusted to strides of [4, 8, 16, 32] This downsampling is accomplished through max pooling, while upsampling is achieved using a single transposed convolution layer. The model is trained for 36 epochs using the AdamW optimizer with learning rate of $10^{-4}$, 0.05 weight decay and 16 batch size. In table 3,we condunct experiments on XCiT-S12 models using 16×16 patches with the activations $\frac{1}{14}x^{3}$, $\frac{1}{14}x$ and softmax. We found we couldn’t train with the activations $x^{3}$ and $x$ on this task well so only report the others.
Activation $AP^{b}$ $AP^{b}_{50}$ $AP^{b}_{75}$ $AP^{m}$ $AP^{m}_{50}$ $AP^{m}_{75}$
softmax 44.9 66.1 48.9 40.1 63.1 42.8
$\frac{x^{3}}{14}$ 44.8 66.3 49 40.1 63.1 42.9
$\frac{x}{14}$ 44.8 66.2 49 40.1 63.1 42.8
Table 3: COCO object detection and instance segmentation performance on the mini-val set. All backbones are pretrained on ImageNet-1k, and use Mask R-CNN model. $AP^{b}$: Average Precision for bounding box predictions, $AP^{b}_{50/75}$: Average Precision at an Intersection over Union (IoU) threshold of 0.50/0.75 for bounding box predictions, $AP^{m}$: Average Precision for mask predictions, $AP^{m}_{50/75}$: Average Precision at an Intersection over Union (IoU) threshold of 0.50/0.75 for mask predictions
5.3 Natural language processing(NLP)
To assess the effectiveness of our approach on NLP tasks, we trained models on five benchmarks from the Long Range Arena (LRA) suite Tay et al. (2020): ListOps, Text Classification, Retrieval, Image Classification, and Pathfinder. We evaluated the activations $\frac{x^{3}}{14}$ and $\frac{x}{14}$ against softmax, finding that $x^{3}$ and $x$ did not train effectively on their own, so only results for these scaled activations and softmax are presented. Our implementation followed the guidelines from Xiong et al. (2021). The results are summarized in table 4.
Activation ListOps Text Retrieval Image Pathfinder
softmax 37.1 63.8 79.8 39.9 72.9
$\frac{x^{3}}{14}$ 37.5 63.3 80.9 37.2 68.7
$\frac{x}{14}$ 34.3 62.9 81.5 39.0 69.1
Table 4: Comparsions of transformer models with different activation functions on NLP tasks. We report the accuracy (%) on LRA benchmarks.
6 Limitations
While our work introduces novel activations that challenge the conventional softmax approach, there are some limitations to address. Our theoretical framework is primarily designed for dot-product self-attention and may not immediately extend to other attention mechanisms, although our empirical results showed competitive performance against softmax across different architectures. Additionally, we observed that while our activations performed well on vision tasks, their performance was less consistent on NLP tasks, suggesting that a more refined theoretical approach may be needed for these applications.
7 Conclusion
This work challenges the traditional view that transformer activations for attention must produce sparse probability distributions. We introduced a theoretical framework analyzing the Frobenius norm of the self-attention matrix, which suggests key scaling properties for activations in attention mechanisms. We proved that specific polynomial activations, which behave very differently from softmax, satisfy these properties. Through extensive experiments across vision and NLP tasks, we demonstrated that these alternative activations not only compete with but sometimes outperform softmax, offering a fresh perspective on attention mechanisms in transformers.
Appendix A Appendix
A.1 Theoretical analysis
A.1.1 Proofs for theorems in section 4.1
In this section we give the proof of theorem 4.1.
Proof of theorem 4.1.
We will start by proving the first inequality in theorem 4.1. Given a matrix $A=(a_{ij})\in\mathbb{R}^{N\times N}$ we have that
$\mathbf{softmax}\bigg{(}\begin{bmatrix}a_{11}&\cdots&a_{1n}\\ \vdots&\vdots&\vdots\\ a_{n1}&\cdots&a_{nn}\end{bmatrix}\bigg{)}=\begin{bmatrix}\frac{e^{a_{11}}}{\sum_{j=1}^{n}e^{a_{1j}}}&\cdots&\frac{e^{a_{1n}}}{\sum_{j=1}^{n}e^{a_{1j}}}\\ \vdots&\vdots&\vdots\\ \frac{e^{a_{n1}}}{\sum_{j=1}^{n}e^{a_{nj}}}&\cdots&\frac{e^{a_{nn}}}{\sum_{j=1}^{n}e^{a_{nj}}}.\end{bmatrix}$ (A.1)
By definition of the Frobenius norm we then see that
$||\mathbf{softmax}(A)||_{F}^{2}$ $=\bigg{(}\frac{1}{\sum_{j=1}^{n}e^{a_{1j}}}\bigg{)}^{2}(e^{2a_{11}}+\cdots e^{2a_{11}})+\cdots+\bigg{(}\frac{1}{\sum_{j=1}^{n}e^{a_{Nj}}}\bigg{)}^{2}(e^{2a_{N1}}+\cdots e^{2a_{NN}})$ (A.2)
$\leq\bigg{[}\bigg{(}\frac{1}{\sum_{j=1}^{n}e^{a_{1j}}}\bigg{)}(e^{a_{11}}+\cdots e^{a_{11}})\bigg{]}^{2}+\cdots+\bigg{[}\bigg{(}\frac{1}{\sum_{j=1}^{n}e^{a_{Nj}}}\bigg{)}(e^{a_{N1}}+\cdots e^{a_{NN}})\bigg{]}^{2}$ (A.3)
$=1+\cdots+1$ (A.4)
$=N$ (A.5)
where the second inequality uses the fact that for non-negative numbers $a$ and $b$ we always have that $a^{2}+b^{2}\leq(a+b)^{2}$.
It then immediately follows that $||\mathbf{softmax}(A)||_{F}\leq\sqrt{N}$ and this proves the first inequality in the statement of theorem 4.1.
We move on to prove the second inequality in the statement of theorem 4.1. For this, let us write each entry of the matrix on the right of equation A.1 as follows:
$F_{kl}=\frac{e^{a_{kl}}}{\sum_{j=1}^{N}e^{a_{kj}}}.$ (A.6)
By applying the chain rule we then have the following derivative formulas
$\frac{\partial}{\partial x_{ij}}F_{ij}$ $=F_{ij}-F_{ij}^{2}$ (A.7)
$\frac{\partial}{\partial x_{ik}}F_{ij}$ $=-F_{ij}F_{ik}\text{ for any }k\neq j$ (A.8)
$\frac{\partial}{\partial x_{kl}}F_{ij}$ $=0\text{ for any }k\neq i\text{ and }l\neq j.$ (A.9)
We can then express the gradient as
$\nabla\mathbf{softmax}(A)=\begin{bmatrix}\nabla F_{11}\\ \vdots\\ \nabla F_{1N}\\ \nabla F_{21}\\ \vdots\\ \vdots\\ \nabla F_{NN}\end{bmatrix}$ (A.10)
where
$\nabla F_{ij}=\begin{bmatrix}-F_{ij}F_{i1}^{2}&-F_{ij}F_{i2}&\cdots&F_{ij}-F_{ij}^{2}&\cdots&F_{ij}F_{iN}.\end{bmatrix}$ (A.11)
From these computations we see that
$||\nabla\mathbf{softmax}(A)||_{F}^{2}=||\nabla F_{11}||_{F}^{2}+\cdots+||\nabla F_{NN}||_{F}^{2}.$ (A.12)
We will proceed by bounding each collection $||\nabla F_{i1}||_{F}^{2}+\cdots+||\nabla F_{1N}||_{F}^{2}$ separately then add up all the bounds. We have
$||\nabla F_{i1}||_{F}^{2}+\cdots+||\nabla F_{1N}||_{F}^{2}$ $=|F_{i1}-F_{i1}^{2}|^{2}+|F_{i1}F_{i2}|^{2}+\cdots+|F_{i1}F_{iN}|^{2}$ (A.13)
$\hskip 14.22636pt+|F_{i2}F_{i1}|^{2}+|F_{i2}-F_{i2}^{2}|^{2}+\cdots+|F_{i2}F_{iN}|^{2}$ (A.14)
[…]
$\hskip 14.22636pt+|F_{iN}F_{i1}|^{2}+|F_{iN}F_{i2}|^{2}+\cdots+|F_{iN}-F_{iN}^{2}|^{2}$ (A.16)
$\leq(F_{i1})^{2}(|1-F_{i1}|+|F_{i2}|+\cdots+|F_{iN}|)^{2}$ (A.17)
$\hskip 14.22636pt(F_{i2})^{2}(|F_{i1}|+|1-F_{i2}|+\cdots+|F_{iN}|)^{2}$ (A.18)
[…]
$\hskip 14.22636pt+(F_{iN})^{2}(|F_{i1}|+|F_{i2}|+\cdots+|1-F_{iN}|)^{2}.$ (A.20)
We then observe that since $F_{i1}+\cdots+F_{iN}=1$ we have that $1-F_{ij}=2(F_{i1}+\cdots+\widehat{F_{ij}}+\cdots+F_{iN})$ where $\widehat{F_{ij}}$ means we don’t include $F_{ij}$ in the sum. This means we get the bound
$||\nabla F_{i1}||_{F}^{2}+\cdots+||\nabla F_{1N}||_{F}^{2}$ $\leq 4F_{i1}^{2}(\widehat{F_{i1}}+F_{i2}+\cdots+F_{iN})$ (A.21)
[…]
$\hskip 14.22636pt+4F_{iN}^{2}(F_{i1}+F_{i2}+\cdots+\widehat{F_{iN}})$ (A.23)
$\leq 4(F_{i1}^{2}+\cdots F_{iN}^{2})$ (A.24)
$=4.$ (A.25)
Putting all the bounds together for each of the terms $N$ terms $||\nabla F_{i1}||_{F}^{2}+\cdots+||\nabla F_{1N}||_{F}^{2}$ we get
$||\nabla\mathbf{softmax}(A)||_{F}^{2}\leq 4N$ (A.26)
and this implies
$||\nabla\mathbf{softmax}(A)||_{F}\leq 2\sqrt{N}.$ (A.27)
This finishes the proof of theorem 4.1.
A.1.2 Proofs for theorems section 4.2
In this section we will give the proof of theorems 4.2 and 4.4.
Proof of theorem 4.2.
We will split the matrix product $XQK^{T}X^{T}$ and think of it as the product of two matrices. Suppose $\mathbf{A}\in\mathbb{R}^{N\times D}\sim\mathcal{N}(0,\sigma_{1}^{2})$, $\mathbf{B}\in\mathbb{R}^{D\times N}\sim\mathcal{N}(0,\sigma_{2}^{2})$ and $\mathbf{C}=\mathbf{AB}$. Each element in the matrix $\mathbf{C}$ can be written as a product of a row of $\mathbf{A}$ with a column of $\mathbf{B}$. Since expectation is linear, we need to compute the expectation of each of these elements. We do the case of the entry $c_{11}$ which is the entry in $\mathbf{C}$ in the first row and first column. For the $p=1$ case we can then compute
$\mathop{\mathbb{E}}(c_{11}^{2})$ $=\mathop{\mathbb{E}}((\sum_{i=1}^{D}a_{1i}b_{i1})^{2})$ (A.28)
$=\mathop{\mathbb{E}}(\sum_{i=1}^{D}a_{1i}^{2}b_{i1}^{2}+\sum_{i=1}^{D}\sum_{j=1,j\neq i}^{D}a_{1i}b_{i1}a_{1j}b_{j1})$
$=\sum_{i=1}^{D}\mathop{\mathbb{E}}(a_{1i}^{2})\mathop{\mathbb{E}}(b_{i1}^{2})+\sum_{i=1}^{D}\sum_{j=1,j\neq i}^{D}\mathop{\mathbb{E}}(a_{1i})\mathop{\mathbb{E}}(b_{i1})\mathop{\mathbb{E}}(a_{1j})\mathop{\mathbb{E}}(b_{j1})$
$=D\sigma_{1}^{2}\sigma_{2}^{2}+0.$
The Frobenius norm of the matrix $\mathbf{C}$ is just the sum of these values for all $N^{2}$ elements and this proves the $p=1$ case.
For the case that $p>1$ we proceed in a similar way. The key observation is that odd powers, in the matrix expansion, will have expectaion 00, so we need only consider the even powers. Therefore, suppose $\mathbf{C}=(\mathbf{A}\mathbf{B})^{p}$. We will compute the expectation of the first entry $c_{1}1\in\mathbf{C}$:
$\mathop{\mathbb{E}}(c_{11}^{2})$ $=\mathop{\mathbb{E}}((\sum_{i=1}^{D}a_{1i}b_{i1})^{2p})$ (A.29)
$=\mathop{\mathbb{E}}(\sum_{i=1}^{D}a_{1i}^{2p}b_{i1}^{2p}+\sum_{i=1}^{D}\sum_{j=1,j\neq i}^{D}a_{1i}^{2p-2}b_{i1}^{2p-2}a_{1j}^{2}b_{j1}^{2}+\cdots).$
Note that the first term only has a count of $D$ and the second term has a count of $D(D-1)$. Thus, we only need to consider the $\mathcal{O}(D^{p})$ term where all the components have a power of 2. The count is similar to choosing $p$ items from $D$,
$\mathop{\mathbb{E}}(c_{11}^{2})$ $\approx\mathop{\mathbb{E}}(\sum_{\{i_{1},\dots,i_{p}\}\in\{1,\dots,D\}}\prod_{k=1}^{p}a_{1,i_{k}}^{2}b_{i_{k},1}^{2})$ (A.30)
$=\left(\begin{array}[]{c}D\\ p\end{array}\right)\frac{2p!}{2^{p}}\sigma_{1}^{2p}\sigma_{2}^{2p}$
$=\frac{D!}{(D-p)!}\frac{2p!}{p!2^{p}}\sigma_{1}^{2p}\sigma_{2}^{2p}$
$=\frac{D!}{(D-p)!}\frac{2p!}{2p!!}\sigma_{1}^{2p}\sigma_{2}^{2p}$
$=\frac{D!}{(D-p)!}(2p-1)!!\sigma_{1}^{2p}\sigma_{2}^{2p}.$
$\frac{D!}{(D-p)!}$ can always be bounded above by $D^{p}$, so the expectation can be upper bounded by $D^{p}(2p-1)!!\sigma_{1}^{2p}\sigma_{2}^{2p}$ and thus we get a quantity of the form $\mathcal{O}(N)$.
Proof of theorem 4.4.
We will do the $p=1$ case first. We proceed similar to the proof of Theorem 4.2.
$\mathop{\mathbb{E}}(\|\frac{\partial XQK^{T}X^{T}}{\partial Q}\|^{2}_{F})$ $=\sum_{i=1}^{N}\sum_{j=1}^{N}\mathop{\mathbb{E}}(|\frac{\partial x_{i}^{T}QK^{T}x_{j}}{\partial Q}|^{2}_{F})$ (A.31)
$=\sum_{i=1}^{N}\sum_{j=1}^{N}\mathop{\mathbb{E}}(\|x_{i}x_{j}^{T}K\|^{2}_{F})$ (A.32)
$=\sum_{i=1}^{N}\sum_{j=1}^{N}\mathop{\mathbb{E}}(\sum_{k=1}^{D}\sum_{l=1}^{d}(x_{ik}\sum_{m=1}^{D}x_{jm}k_{ml})^{2})$ (A.33)
$=\sum_{i=1}^{N}\sum_{j=1}^{N}\mathop{\mathbb{E}}(\sum_{k=1}^{D}\sum_{l=1}^{d}x_{ik}^{2}(\sum_{m=1}^{D}x_{jm}k_{ml})^{2})$ (A.34)
$=\sum_{i=1}^{N}\sum_{j=1}^{N}\mathop{\mathbb{E}}(\sum_{k=1}^{D}\sum_{l=1}^{d}x_{ik}^{2}(\sum_{m=1}^{D}x_{jm}^{2}k_{ml}^{2}+\sum_{m=1}^{D}\sum_{n=1,n\neq m}^{D}x_{jm}k_{ml}x_{jn}k_{nl}))$ (A.35)
$=\sum_{i=1}^{N}\sum_{j=1}^{N}\sum_{k=1}^{D}\sum_{l=1}^{d}(\sum_{m=1}^{D}\mathop{\mathbb{E}}(x_{ik}^{2}x_{jm}^{2}k_{ml}^{2})+\sum_{m=1}^{D}\sum_{n=1,n\neq m}^{D}\mathop{\mathbb{E}}(x_{ik}^{2}x_{jm}k_{ml}x_{jn}k_{nl}))$ (A.36)
$=\sum_{i=1}^{N}\sum_{j=1}^{N}\sum_{k=1}^{D}\sum_{l=1}^{d}(\sum_{m=1}^{D}\mathop{\mathbb{E}}(x_{ik}^{2}x_{jm}^{2}k_{ml}^{2})+0)$ (A.37)
$=\sum_{i=1}^{N}\sum_{j=1}^{N}\sum_{k=1}^{D}\sum_{l=1}^{d}\sum_{m=1}^{D}\mathop{\mathbb{E}}(x_{ik}^{2}x_{jm}^{2}k_{ml}^{2})$ (A.38)
$=\sum_{i=1}^{N}\sum_{k=1}^{D}\sum_{l=1}^{d}\mathop{\mathbb{E}}(x_{ik}^{2}x_{ik}^{2}k_{kl}^{2})+\sum_{i=1}^{N}\sum_{j=1,j\neq i}^{N}\sum_{k=1}^{D}\sum_{l=1}^{d}\sum_{m=1,m\neq k}^{D}\mathop{\mathbb{E}}(x_{ik}^{2}x_{jm}^{2}k_{ml}^{2})$ (A.39)
$=NDd3\sigma_{x}^{4}\sigma_{w}^{2}+N(N-1)D(D-1)d\sigma_{x}^{4}\sigma_{w}^{2}$ (A.40)
$\approx N^{2}D^{2}d\sigma_{x}^{4}\sigma_{w}^{2}.$ (A.41)
When $p>1$ we can proceed in a similar way.
$\mathop{\mathbb{E}}(\|\frac{\partial(XQK^{T}X^{T})^{p}}{\partial Q}\|^{2}_{F})$ $=\sum_{i=1}^{N}\sum_{j=1}^{N}\mathop{\mathbb{E}}(\|\frac{(\partial x_{i}^{T}QK^{T}x_{j})^{p}}{\partial Q}\|^{2}_{F})$ (A.42)
$=\sum_{i=1}^{N}\sum_{j=1}^{N}\mathop{\mathbb{E}}(\|p(x_{i}^{T}QK^{T}x_{j})^{p-1}\frac{\partial x_{i}^{T}QK^{T}x_{j}}{\partial Q}\|^{2}_{F})$ (A.43)
$=\sum_{i=1}^{N}\sum_{j=1}^{N}\mathop{\mathbb{E}}(\|p(x_{i}^{T}QK^{T}x_{j})^{p-1}x_{i}x_{j}^{T}K\|^{2}_{F})$ (A.44)
$=\sum_{i=1}^{N}\sum_{j=1}^{N}\mathop{\mathbb{E}}(p^{2}(x_{i}^{T}QK^{T}x_{j})^{2p-2}\sum_{k=1}^{D}\sum_{l=1}^{d}(x_{ik}\sum_{m=1}^{D}x_{jm}k_{ml})^{2}).$ (A.45)
We know that
$(x_{i}^{T}QK^{T}x_{j})^{2p-2}$ $=(\sum_{l=1}^{d}((\sum_{k=1}^{D}x_{ik}q_{kl})\cdot(\sum_{m=1}^{D}x_{jm}k_{ml})))^{2p-2}$ (A.46)
$=(\sum_{l=1}^{d}\sum_{k=1}^{D}\sum_{m=1}^{D}x_{ik}q_{kl}x_{jm}k_{ml})^{2p-2}$ (A.47)
$=(\sum_{k=1}^{D}\sum_{m=1}^{D}x_{ik}x_{jm}\sum_{l=1}^{d}q_{kl}k_{ml})^{2p-2}$ (A.48)
$=(\sum_{k=1}^{D}\sum_{m=1}^{D}x_{ik}x_{jm}a_{km})^{2p-2},$ (A.49)
where $a_{km}=\sum_{l=1}^{d}q_{kl}k_{ml}$. Let $z_{ij}=\sum_{k=1}^{D}\sum_{m=1}^{D}x_{ik}x_{jm}a_{km}$ Thus we have
$\mathop{\mathbb{E}}(\|\frac{\partial(XQK^{T}X^{T})^{p}}{\partial Q}\|^{2}_{F})$ $=$ $p^{2}\sum_{i=1}^{N}\sum_{j=1}^{N}\mathop{\mathbb{E}}(z_{ij}^{2p-2}\sum_{k=1}^{D}\sum_{l=1}^{d}(x_{ik}\sum_{m=1}^{D}x_{jm}k_{ml})^{2})$ (A.50)
$=$ $\sum_{i=1}^{N}\sum_{j=1}^{N}\sum_{k=1}^{D}\sum_{l=1}^{d}(\sum_{m=1}^{D}\mathop{\mathbb{E}}(z_{ij}^{2p-2}x_{ik}^{2}x_{jm}^{2}k_{ml}^{2})$ (A.51)
$+\sum_{m=1}^{D}\sum_{n=1,n\neq m}^{D}\mathop{\mathbb{E}}(z_{ij}^{2p-2}x_{ik}^{2}x_{jm}k_{ml}x_{jn}k_{nl}))$ (A.52)
$=$ $\sum_{i=1}^{N}\sum_{j=1}^{N}\sum_{k=1}^{D}\sum_{l=1}^{d}(\sum_{m=1}^{D}\mathop{\mathbb{E}}(z_{ij}^{2p-2}x_{ik}^{2}x_{jm}^{2}k_{ml}^{2})$ (A.53)
$+\sum_{m=1}^{D}\sum_{n=1,n\neq m}^{D}\mathop{\mathbb{E}}(z_{ij}^{2p-3}x_{ik}^{2}x_{jm}^{2}k_{ml}^{2}x_{jn}^{2}k_{nl}^{2}))$ (A.54)
$\approx$ $N^{2}Dd(D(D^{2p-2}d^{p-1}(2p-3)!!\sigma_{x}^{4p}\sigma_{w}^{4p-2})+0$ (A.55)
$=$ $N^{2}D^{2p}d^{p}(2p-3)!!\sigma_{x}^{4p}\sigma_{w}^{4p-2}$ (A.56)
showing that we can bound the gradient by a quantity of the form $\mathcal{O}(N)$ and the proof is complete.
A.2 Experiments
A.2.1 Ablations
In this section we carry out ablations on the experiments we did in the main paper.
Scale ablations:
The theory we developed in section 4.2 suggested that we needed to scale our polynomial activations by $\frac{1}{\sqrt{N}}$ to obtain a complexity bound of $\mathcal{O}(\sqrt{N})$. In general, we could also scale the polynomial activations by $\mathcal{O}(\frac{1}{\sqrt{N}})$ to see if we can get a better accuracr. Figure 10 carries out an ablation on both the ViT-Base architecture and the ViT-Small architecture to see how different scales affect the accuracy for the $x^{3}$ activation. We found that in general a scale from $\frac{1}{8}$ to $\frac{1}{25}$ seemed to perform very well.
Figure 10: We show how the top-1% accuracy changes as the scale of the activation $x^{3}$ for a ViT-Base (left) and ViT-Small (right) architecture. The x-axis plots the denominator of the scale for easier readability. In other words as we go to the right of the x-axis the scale used on the activation $x^{3}$ is getting smaller.
Activation ablations:
Our theory primarily compared the activations $\frac{1}{\sqrt{N}}x^{3}$ and $\frac{1}{\sqrt{N}}x$ with softmax where $N$ was the sequence length of the input. In this section we carry out further experimental comparisons of our activations with other activations. We compare with a variety of activations used in the literature as well as the exponential function $Exp=e^{-x}$ which can be thought of as softmax without the row normalization scaling. Table 5 shows the results of the experiments.
Activation ViT-Base ViT-small
softmax 79.6 80.2
$\frac{x^{3}}{14}$ 79.6 80.5
$\frac{x}{14}$ 76.9 77.8
$\frac{x^{2}}{14}$ 78.7 79.9
$ReLU$ 77.3 77.5
$ELU$ 78.2 78.5
$GELU$ 78.2 78.4
$Tanh$ 77.1 77.3
$Exp$ 77.9 78.0
Table 5: Top-1% accuracy with various activations on a ViT-Base and ViT-Small architecture.
A.2.2 Frobenius norm computations
In section 5.1 we showed plots of the Frobenius norm of the self-attention matrix and for the Jacobian of the self-attention matrix for different scalings of the $x^{3}$ and $x$ activations along with softmax. This was done for a ViT-Tiny architecture on the Tiny-ImageNet dataset. Figure 11 shows the plots of the Frobenius norm of the self-attention matrix for the Tiny-ViT architecture, during training, for all layers averaged over the heads within each layer. Figure 12 shows the Frobenius norm of the Jacobian of the self-attention matrix during training for each layer, averaged over the total number of heads within each layer.
Figures 13 and 14 show similar results for the ViT-Small architecture.
Figure 11: Frobenius norm of self-attention matrix for different scaled activations on ViT-Tiny during training (zoom in for better viewing).
Figure 12: Frobenius norm of Jacobian of self-attention matrix for different scaled activations on ViT-Tiny during training (zoom in for better viewing).
Figure 13: Frobenius norm of self-attention matrix for different scaled activations on ViT-Small during training (zoom in for better viewing).
Figure 14: Frobenius norm of Jacobian of self-attention matrix for different scaled activations on ViT-Small during training (zoom in for better viewing).

### v2

@@
-Rethinking Softmax: Self-Attention with Polynomial Activations
+Rethinking Attention: Polynomial Alternatives to Softmax in Transformers
 Abstract
-This paper challenges the conventional belief that softmax attention in transformers is effective primarily because it generates a probability distribution for attention allocation.
-Instead, we theoretically show that its success lies in its ability to implicitly regularize the Frobenius norm of the attention matrix during training.
-We then explore alternative activations that regularize the Frobenius norm of the attention matrix, demonstrating that certain polynomial activations can achieve this effect, making them suitable for attention-based architectures.
-Empirical results indicate these activations perform comparably or better than softmax across various computer vision and language tasks, suggesting new possibilities for attention mechanisms beyond softmax.
+This paper questions whether the strong performance of softmax attention in transformers stems from producing a probability distribution over inputs.
+Instead, we argue that softmax’s effectiveness lies in its implicit regularization of the Frobenius norm of the attention matrix, which stabilizes training.
+Motivated by this, we explore alternative activations, specifically polynomials, that achieve a similar regularization effect.
+Our theoretical analysis shows that certain polynomials can serve as effective substitutes for softmax, achieving strong performance across transformer applications despite violating softmax’s typical properties of positivity, normalization, and sparsity.
+Extensive experiments support these findings, offering a new perspective on attention mechanisms.
 1 Introduction
-Transformer architectures (Vaswani et al., 2017) have become the state-of-the-art model architecture for diverse areas such as natural language processing (NLP) (Vaswani et al., 2017; Devlin et al., 2018; Zhuang et al., 2021; Zhen et al., 2022), computer vision (Dosovitskiy et al., 2020; Carion et al., 2020; Liu et al., 2021; Touvron et al., 2021), and robotics (Fu et al., 2024; Maiti et al., 2023; Salzmann et al., 2020).
-A key component in the transformer architecture is the softmax attention block, enabling transformers to evaluate the importance of individual input elements during output generation.
-This feature facilitates an efficient method to attend to diverse input elements throughout training, allowing transformers to effectively capture spatial dependencies within sequential data.
-Unlike traditional recurrent neural networks (RNNs) and convolutional neural networks (CNNs), transformers exhibit the ability to scale to large datasets without a significant degradation in performance.
-This characteristic has made them an ideal architecture for handling large-scale machine learning tasks.
-Softmax is widely recognized for its effectiveness in attention mechanisms due to its ability to produce an attention matrix that meets three key conditions (Vaswani et al., 2017; Dosovitskiy et al., 2020; Zhen et al., 2022; AUEB et al., 2016): (i) non-negativity, (ii) rows that are normalized to sum to 1 and (iii) sparsity.
-The general consensus is that non-negativity guarantees that attention weights remain positive, facilitating the model in assigning significance to various input elements.
-The normalization constraint ensures that the attention weights for all input elements collectively sum to $1$, rendering the weights interpretable as probabilities.
-Additionally, sparsity aids the model in focusing on a select few key elements, thereby enhancing efficiency.
-It has been argued that these properties are crucial for enabling the attention mechanism to attend to pertinent segments of the input sequence while efficiently filtering out irrelevant details.
-However, this approach to attention has become somewhat axiomatic as it is mostly motivated by empirical results with little theoretical foundation.
-Despite the exploration of alternative activations in several studies (Shen et al., 2023; Fang et al., 2022; Correia et al., 2019), softmax attention continues to dominate, largely due to its interpretability.
-In this paper, we question this view by proposing that the effectiveness of softmax stems from its implicit regularization of the Frobenius norm of the attention matrix during training, preventing attention weights from becoming excessively large or small.
-We then derive a theoretical framework that produces polynomial activations that deliberately violate one or more of the three conditions mentioned earlier, yet are able to regularize the Frobenius norm of the attention weights during training.
-Our findings demonstrate that such activations can achieve comparable or even superior performance to softmax across various vision and natural language processing (NLP) tasks, even though they seem to violate our understanding of attention.
-We advise the reader that this paper diverges from the usual pursuit of creating cutting-edge transformer architectures for achieving state-of-the-art results on benchmark datasets.
-Instead, our focus is on critically examining softmax attention to determine whether its effectiveness is a result of true interpretability or a more nuanced regularization mechanism.
-By questioning established views, we aim to uncover deeper insights into transformer architectures that could lead to broader applications and improved understanding.
-Nonetheless, we validate our theory on multiple transformer-based tasks, including image classification, segmentation, object detection, and NLP, often achieving results that match or surpass those of softmax attention.
-Our main contributions are:
+Transformer architectures [26] have become the foundation of state-of-the-art models across natural language processing (NLP) [26, 9, 33, 32], computer vision [10, 6, 17, 25], and robotics [12, 18, 19].
+A core component of these models is the softmax attention block, which computes token interactions by evaluating the relative importance of inputs.
+This mechanism enables transformers to effectively model long-range dependencies, a capability that distinguishes them from recurrent neural networks (RNNs) and convolutional neural networks (CNNs), particularly in scaling to large datasets.
+Softmax self-attention satisfies three key properties: (1) non-negativity of attention weights, (2) row-wise normalization to ensure weights sum to one (interpretable as probabilities), and (3) sparsity, promoting focus on a few relevant tokens.
+These properties are commonly believed to be essential for modeling effective attention [4, 32], facilitating interpretability.
+However, this viewpoint has largely been adopted on empirical grounds, with limited theoretical justification.
+Despite investigations into alternative activations [20, 11, 8], softmax remains dominant, primarily due to its empirical performance and interpretability.
+In this paper, we challenge this prevailing assumption and ask:
+{greenbox}
+Do attention mechanisms in transformers require non-negativity, normalization, and sparsity for effective performance?
+We present a new perspective suggesting that softmax’s effectiveness stems not inherently from these properties, but from its implicit regularization of the attention matrix’s Frobenius norm during training.
+Building on this insight, we show that simple polynomial activations, which violate non-negativity, normalization, and sparsity, can nevertheless provide comparable or superior performance to softmax by inducing a similar regularization effect.
+This offers a fundamentally different interpretation of attention, distinct from the conventional probabilistic view of softmax.
+Our contributions are as follows:
   • 1.
-    We question the widely accepted notion that softmax’s effectiveness in attention mechanisms is solely due to its ability to produce normalized sparse attention weights.
-Instead, we theoretically show that softmax has a regularization effect on attention and argue this plays a crucial role in its success.
+    We provide a theoretical analysis showing that softmax attention implicitly regularizes the Frobenius norm of the attention matrix, challenging the assumption that non-negativity, normalization, and sparsity are the primary drivers of its success.
   • 2.
-    We explore activations that deliberately deviate from traditional softmax attention conditions.
-These activations are found to regularize the Frobenius norm of the attention matrix during training, akin to softmax, and demonstrate comparable or superior performance across various vision and NLP tasks.
+    We propose polynomial activations as an alternative to softmax, demonstrating that they can induce similar regularization effects without adhering to the traditional softmax constraints, achieving competitive performance across tasks including image classification, object detection, instance segmentation, text classification, and physics-based modeling.
+By revisiting foundational assumptions, our work offers deeper insights into the mechanics of attention and opens new avenues for designing alternative attention mechanisms.
 2 Related Work
-Several studies have explored alternative activations for attention mechanisms in transformers.
-Shen et al.
-(2023) investigated ReLU activations, finding them to outperform softmax in tasks with long sequences, such as document translation.
-Banerjee et al.
-(2020) examined Taylor series approximations to softmax, which showed superior performance to softmax in image classification.
-Wang et al.
-(2021) proposed periodic alternatives to softmax, designed to provide better gradients for attention mechanisms and achieved better results than softmax on simple networks for image classification.
-Koohpayegani & Pirsiavash (2024) demonstrated that applying $l^{1}$ normalization to linear attention mechanisms can achieve performance comparable to softmax.
-Our work differs from all these in that we identify a clear theoretical relationship between the scale of the Frobenius norm of the self-attention matrix and the input sequence length.
-Using this insight to derive potential acitvations that can perform on par with softmax.
+Attention activations.
+A variety of alternative activations for attention mechanisms have been explored in recent literature.
+[20] showed that ReLU activations outperform softmax in long-sequence tasks, such as document translation.
+Of particular relevance to our work, [29] demonstrated that scaling ReLU by the inverse of sequence length can surpass softmax in certain vision applications, emphasizing the importance of correct activation scaling.
+In this paper, we show that softmax inherently applies such a scale through its normalization, and we derive theoretical principles that motivate polynomial activations with scalings proportional to the square root of the sequence length.
+Other studies have proposed alternatives with varying motivations.
+[5] used Taylor series approximations of softmax, achieving superior performance in image classification.
+[27] introduced periodic activations to improve gradient flow in attention layers.
+[15] showed that $l^{1}$ normalization applied to linear attention mechanisms can yield on par performance to softmax’s on three distinct vision transformers.
+Distinct from these works, our approach establishes a clear theoretical link between the Frobenius norm of the attention matrix and the input sequence length.
+Leveraging this insight, we design polynomial activations that break three canonical properties of softmax—non-negativity, row normalization, and sparsity—while still achieving competitive performance.
+Attention mechanisms.
+Numerous strategies have been proposed to improve the efficiency and scalability of transformers by reducing computational overhead and rethinking attention mechanisms.
+The Data-Efficient Image Transformer (DeiT) [25] leverages distillation tokens to achieve competitive performance without relying on large datasets.
+The Cross-Covariance Image Transformer (XCiT) [3] introduces cross-covariance attention, enabling efficient spatial interactions with reduced complexity.
+The Swin Transformer [17] employs a hierarchical architecture with shifted window-based self-attention to enhance scalability for vision tasks.
+The Nyströmformer [30] approximates full self-attention using the Nyström method, reducing its complexity from quadratic to near-linear.
+Similarly, the MLP-Mixer [24] replaces self-attention entirely with multi-layer perceptrons for spatial and channel mixing.
+In this work, we demonstrate that our polynomial-based attention activations can be integrated into these architectures [25, 3, 17, 30], matching softmax performance while violating its typical properties of non-negativity, normalization, and sparsity.
 3 Preliminaries and Notation
 In this section we outline the definition of a transformer via the transformer block and set the notation of various mathematical quantities we will be using in future sections.
-For more details on transformers the reader can consult Vaswani et al.
-(2017); Dosovitskiy et al.
-(2020).
+For more details on transformers the reader can consult [26, 10].
 Transformer architectures comprise of transformer blocks, defined as follows. A transformer block is a mapping $\mathbf{T}:\mathbb{R}^{N\times D}\rightarrow\mathbb{R}^{N\times D}$ defined as
-$\mathbf{T}(x)=\mathbf{F}(\mathbf{A}(x)+x)$ (3.1)
+$\mathbf{T}(x)=\mathbf{F}(\mathbf{A}(x)+x)$ (1)
 where $\mathbf{F}$ is a feedforward MLP with a residual connection and $\mathbf{A}$ is an attention head.
@@
 The attention head $\mathbf{A}(X)$ is then defined by
-$\mathbf{A}(X)=\phi(\mathcal{S}(q,k))v$ (3.2)
+$\mathbf{A}(X)=\phi(\mathcal{S}(q,k))v$ (2)
 where $\mathcal{S}$ is a similarity transformation and $\phi$ is an activation function.
@@
 This leads to the most common form of the attention head given by
-$\mathbf{A}(X)=\mathbf{softmax}\bigg{(}\frac{qk^{T}}{\sqrt{d}}\bigg{)}v=\mathbf{softmax}\bigg{(}\frac{XQK^{T}X^{T}}{\sqrt{d}}\bigg{)}XV.$ (3.3)
+$\begin{split}\mathbf{A}(X)&=\mathbf{softmax}\bigg{(}\frac{qk^{T}}{\sqrt{d}}\bigg{)}v\\ &=\mathbf{softmax}\bigg{(}\frac{XQK^{T}X^{T}}{\sqrt{d}}\bigg{)}XV.\end{split}$ (3)
 The function $\mathbf{softmax}$ is the matrix softmax map that applies the usual softmax function row-wise:
-$\mathbf{softmax}\bigg{(}\begin{bmatrix}x_{11}&\cdots&x_{1n}\\ \vdots&\vdots&\vdots\\ x_{n1}&\cdots&x_{nn}\end{bmatrix}\bigg{)}=\begin{bmatrix}\frac{e^{x_{11}}}{\sum_{j=1}^{n}e^{x_{1j}}}&\cdots&\frac{e^{x_{1n}}}{\sum_{j=1}^{n}e^{x_{1j}}}\\ \vdots&\vdots&\vdots\\ \frac{e^{x_{n1}}}{\sum_{j=1}^{n}e^{x_{nj}}}&\cdots&\frac{e^{x_{nn}}}{\sum_{j=1}^{n}e^{x_{nj}}}\end{bmatrix}$ (3.4)
-The factor $\frac{1}{\sqrt{d}}$, as explained in Vaswani et al.
-(2017), is a scaling to prevent the gradients of softmax from being too small.
+$\mathbf{softmax}\bigg{(}\begin{bmatrix}x_{11}&\cdots&x_{1n}\\ \vdots&\vdots&\vdots\\ x_{n1}&\cdots&x_{nn}\end{bmatrix}\bigg{)}=\begin{bmatrix}\frac{e^{x_{11}}}{\sum_{j=1}^{n}e^{x_{1j}}}&\cdots&\frac{e^{x_{1n}}}{\sum_{j=1}^{n}e^{x_{1j}}}\\ \vdots&\vdots&\vdots\\ \frac{e^{x_{n1}}}{\sum_{j=1}^{n}e^{x_{nj}}}&\cdots&\frac{e^{x_{nn}}}{\sum_{j=1}^{n}e^{x_{nj}}}\end{bmatrix}$ (4)
+The factor $\frac{1}{\sqrt{d}}$, as explained in [26], is a scaling to prevent the gradients of softmax from being too small.
 For the theoretical analysis in this paper we will only use the dot-product similarity $qk^{T}$ and call the $N\times N$ matrix $softmax(qk^{T})$ the softmax self-attention matrix.
-In the experiments section, Sec.
-5, we will empirically validate our theoretical framework on more general softmax attention blocks.
+In the experiments, Section 6, we will empirically validate our theoretical framework on more general softmax attention blocks used in state of the art transformers such as DeiT [25], Swin Transformer [17] and XciT [30].
 For general transformer architectures, multiple heads $\mathbf{A}_{i}$ for $1\leq i\leq n$ are used.
-Each attention head is defined by equation 3.3 and then all outputs of each attention head are concatenated together before going into the feedforward layer.
-We will need notation for the derivative of the matrix softmax map defined by equation 3.4.
+Each attention head is defined by (3) and then all outputs of each attention head are concatenated together before going into the feedforward layer.
+We will need notation for the derivative of the matrix softmax map defined by (4).
 Given a matrix $A\in\mathbb{R}^{N\times N}$ we can differentiate the matrix map $\mathbf{softmax}$ at $A$ and obtain the gradient linear map $\mathbf{\nabla softmax}(A):\mathbb{R}^{N\times N}\rightarrow\mathbb{R}^{N\times N}$ that is defined by the formula
-$\mathbf{\nabla softmax}(A):=\mathbf{Jsoftmax}(A)^{T}$ (3.5)
+$\mathbf{\nabla softmax}(A):=\mathbf{Jsoftmax}(A)^{T}$ (5)
 where $\mathbf{Jsoftmax}(A)$ is the Jacobian of $\mathbf{softmax}$ at $A$.
@@
 Additionally, we demonstrate that the gradient of the softmax with respect to the self-attention matrix also exhibits a similar degree of regularity.
-While previous work has analyzed the regularity of softmax self-attention through the lens of the Lipschitz constant (Kim et al., 2021; Castin et al., 2023), our theorem offers a novel perspective by directly linking the Frobenius norm regularity to the token length.
+While previous work has analyzed the regularity of softmax self-attention through the lens of the Lipschitz constant [14, 7], our theorem offers a novel perspective by directly linking the Frobenius norm regularity to the token length.
 This provides insights into how self-attention activations should scale with token length to maintain stability during training, especially with gradient descent-based algorithms.
+{bluebox}
 Theorem 4.1.
-Let $\mathbf{softmax}:\mathbb{R}^{N\times N}\rightarrow\mathbb{R}^{N\times N}$ be the matrix softmax map defined by equation 3.4 and let $\mathbf{\nabla softmax}(A):\mathbb{R}^{N\times N}\rightarrow\mathbb{R}^{N\times N}$ denote the gradient of $\mathbf{softmax}$ at $A\in\mathbb{R}^{N\times N}$.
+Let $\mathbf{softmax}:\mathbb{R}^{N\times N}\rightarrow\mathbb{R}^{N\times N}$ be the matrix softmax map defined by (4) and let $\mathbf{\nabla softmax}(A):\mathbb{R}^{N\times N}\rightarrow\mathbb{R}^{N\times N}$ denote the gradient of $\mathbf{softmax}$ at $A\in\mathbb{R}^{N\times N}$.
 We then have the following bounds on the Frobenius norms
-$||\mathbf{softmax}(A)||_{F}$ $\leq\sqrt{N}$ (4.1)
-$||\mathbf{\nabla softmax}(A)||_{F}$ $\leq 2\sqrt{N}.$ (4.2)
+$||\mathbf{softmax}(A)||_{F}$ $\leq\sqrt{N}$ (6)
+$||\mathbf{\nabla softmax}(A)||_{F}$ $\leq 2\sqrt{N}.$ (7)
 The key implication of theorem 4.1 is that during the training of a transformer with softmax self-attention, the Frobenius norm of each softmax self-attention matrix remains bounded by a value that grows as $\mathcal{O}(\sqrt{N})$.
@@
 The proof hinges on the fact that the row normalization inherent in softmax effectively controls the Frobenius norm.
-For a detailed proof see appendix A.1.1.
+For a detailed proof see Section A.1.1.
 4.2 Polynomial activations for self-attention
@@
 Building on this, we now show that by scaling specific polynomial activations, a similar regularization effect on the Frobenius norm can be achieved in expectation, closely replicating the impact of softmax.
+{bluebox}
 Theorem 4.2.
@@
 We have the following expectations of the Frobenius norms of powers of the $N\times N$ matrix $(XQK^{T}X^{T})^{p}$ for $p\geq 1$
-$\mathbb{E}\bigg{|}\bigg{|}\bigg{(}\frac{XQK^{T}X^{T}}{\sqrt{d}}\bigg{)}^{p}\bigg{|}\bigg{|}_{F}\leq\mathcal{O}(N)$ (4.3)
+$\mathbb{E}\bigg{|}\bigg{|}\bigg{(}\frac{XQK^{T}X^{T}}{\sqrt{d}}\bigg{)}^{p}\bigg{|}\bigg{|}_{F}\leq\mathcal{O}(N)$ (8)
 By scaling such an activation by $\frac{1}{\sqrt{N}}$ we can obtain a $\mathcal{O}(\sqrt{N})$ bound.
+{orangebox}
 Corollary 4.3.
 Assume the same conditions as in theorem 4.2. Then
-$\mathbb{E}\bigg{|}\bigg{|}\frac{1}{\sqrt{N}}\bigg{(}\frac{XQK^{T}X^{T}}{\sqrt{d}}\bigg{)}^{p}\bigg{|}\bigg{|}_{F}\leq\mathcal{O}(\sqrt{N}).$ (4.4)
+$\mathbb{E}\bigg{|}\bigg{|}\frac{1}{\sqrt{N}}\bigg{(}\frac{XQK^{T}X^{T}}{\sqrt{d}}\bigg{)}^{p}\bigg{|}\bigg{|}_{F}\leq\mathcal{O}(\sqrt{N}).$ (9)
 Corollary 4.3 establishes that activations of the form $\phi(x):=\frac{1}{\sqrt{N}}x^{p}$ provide a level of regularization, in expectation, similar to that of softmax when applied to the self-attention matrix.
@@
 The next property we want to prove is one similar to the gradient bound obtained in Theorem 4.1.
-Since the self-attention matrix has parameters given by the queries $Q$ and keys $K$ (Vaswani et al., 2017), this implies that during the training of a transformer the $Q$ and $K$ matrices are the only aspects of the self-attention matrix that get updated.
-Therefore, we compute a regularity result with respect to the $Q$ and $K$ derivatives.
+Since the self-attention matrix has parameters given by the queries $Q$ and keys $K$ [26], this implies that during the training of a transformer the $Q$ and $K$ matrices are the only aspects of the self-attention matrix that get updated.
+Therefore, we compute a derivative bound with respect to the $Q$ and $K$ derivatives.
+{bluebox}
 Theorem 4.4.
@@
 Then the expectation of the of the derivative of the matrix $\frac{(XQK^{T}X^{T})^{p}}{\sqrt{d}}$ w.r.t the $Q$ parameter matrix for $p\geq 1$ is given by
-$\mathbb{E}\bigg{|}\bigg{|}\frac{\partial}{\partial Q}\bigg{(}\frac{(XQK^{T}X^{T})^{p}}{\sqrt{d}}\bigg{)}\bigg{|}\bigg{|}\leq\mathcal{O}(N)$ (4.5)
-The above theorem then suggests that if we scale the polynomial $x\rightarrow x^{p}$ by $\frac{1}{\sqrt{N}}$ the $Q$ derivative will have $\mathcal{O}(\sqrt{N})$ growth.
+$\mathbb{E}\bigg{|}\bigg{|}\frac{\partial}{\partial Q}\bigg{(}\frac{(XQK^{T}X^{T})^{p}}{\sqrt{d}}\bigg{)}\bigg{|}\bigg{|}\leq\mathcal{O}(N)$ (10)
+Figure 1: Training ViT-Tiny with the activation $\phi(x)=x^{3}$ with different sequence lengths and different scales.
+As the sequence length gets larger, the $k$ scale (x-axis) needed to obtain good accuracy when using $\frac{1}{k}x^{3}$ as an activation increases validating the theory from section 4.2.
+{orangebox}
 Corollary 4.5.
 Assume the same condition as in theorem 4.4. Then
-$\mathbb{E}\bigg{|}\bigg{|}\frac{1}{\sqrt{N}}\frac{\partial}{\partial Q}\bigg{(}\frac{(XQK^{T}X^{T})^{p}}{\sqrt{d}}\bigg{)}\bigg{|}\bigg{|}\leq\mathcal{O}(\sqrt{N}).$ (4.6)
+$\mathbb{E}\bigg{|}\bigg{|}\frac{1}{\sqrt{N}}\frac{\partial}{\partial Q}\bigg{(}\frac{(XQK^{T}X^{T})^{p}}{\sqrt{d}}\bigg{)}\bigg{|}\bigg{|}\leq\mathcal{O}(\sqrt{N}).$ (11)
 An analogous estimate holds for derivatives with respect to the $K$ matrix. The proof of theorem 4.4 can be found in appendix A.1.2.
-Corollaries 4.3 and 4.5 suggest that polynomial activations of the form $\phi(x)=\frac{1}{\sqrt{N}}x^{p}$, with $p>0$, can achieve performance comparable to softmax when applied to self-attention matrices.
-In section 5, we empirically compare these activations to softmax and observe that they outperform softmax on a variety of transformer tasks.
-We focus on $p=1$ and $p=3$ as these polynomials clearly violate key aspects of softmax based attention, such as normalized rows, positivity, and sparsity.
-For larger values of $p$, performance declines due to the functions $\phi(x)=\frac{1}{\sqrt{N}}x^{p}$ having smaller gradients around 00 when $p$ is large, causing difficulties in training.
+Remark 4.6.
+Corollary 4.3 and Corollary 4.5 suggest that polynomial activations of the form $\phi(x)=\frac{1}{\sqrt{N}}x^{p}$, with $p>0$, can achieve performance comparable to softmax when applied to self-attention matrices.
+We point out that both corollaries rested on the assumption that $X$, $Q$ and $K$ are i.i.d random variables.
+In general, this is only true at initialization when training a transformer [2].
+Although this is a limitation in the theory the experiments in Section 6 show that this insight can be used effectively to develop new attention blocks that perform comparable to softmax yet violate the three conditions of positivity, normalized rows summing to $1$ and sparsity showing that attention blocks do not need to be modeled as a probability distribution.
+5 Testing the theory
+In this section we test the theory developed in Section 4.2 on small vision transformers.
+We will consider the activation $\phi(x)=\frac{1}{k}x^{3}$, where $k>0$ is a fixed scale, as this activation clearly violates the three key conditions of softmax based attention; positivity, normalization and sparsity.
+We found that polynomials $\frac{1}{k}x^{p}$ for $p>3$ did not perform well during training as they witnessed a gradient vanishing problem due to the fact that the function $x^{p}$ for $p$ large have very small values around 00.
+The first experiment we conducted was to test how the Top-$1\%$ accuracy changes for a ViT-Tiny vision transformer [22], trained on Tiny-Imagenet dataset [21], as we change the sequence length $N$ of the input and the scale predicted in corollaries 4.3 and 4.5 when using the activation $\frac{1}{k}x^{3}$.
+The standard ViT-Tiny model comprises $12$ transformer layers, each equipped with $3$ attention heads, with each head having dimension $64$.
+We considered four different input sequence lengths $N$ of sizes $256$, $64$, $16$ and $8$.
+For each such sequence length, we ran a ViT-Tiny architecture with the activation $\frac{1}{k}x^{3}$ where $k$ ranged from roughly $10^{-3}$ to $10^{3}$.
+According to the theory developed in section 4, the Frobenius norm of $\frac{1}{\sqrt{N}}x^{3}$ scales according to $\mathcal{O}(\sqrt{N})$.
+Thus the best accuracy should occur when $k=\mathcal{O}(\sqrt{N})$ and should degrade for other values due to training instability.
+Figure 1 shows the results of the experiment, we note that the $x-axis$ plots the values of $k$.
+We see that as the sequence length increases the factor of $k$ needs to increase so that the activation $\frac{1}{k}x^{3}$ performs well on the ViT-Tiny architecture as predicted by the theory developed in corollaries 4.3 and 4.5.
+In a second experiment we decided to compare the performance of the original ViT-Tiny architecture, that uses an input sequence length of $256$, with a softmax activation and a polynomial activation on the Tiny-ImageNet dataset.
+In Corollary 4.3 and Corollary 4.5 it was pointed out that the scaling of the polynomial is important to keep the Frobenius norm of the polynomial attention matrix from becoming too large.
+The results of those corollaries suggested that a scale to use is $\frac{1}{\sqrt{N}}$, $N$ being the sequence length, which in this case is $\frac{1}{\sqrt{256}}=\frac{1}{16}$.
+We therefore decided to to compare the three activations softmax, $x^{3}$ and $\frac{1}{16}x^{3}$.
+Each was trained for 200 epochs using the AdamW optimizer.
+To begin with we computed the Frobenius norm of the attention matrix throughout training averaged over all the heads in layers 2, 7 and 12 of the ViT-Tiny architecture when trained on the Tiny-ImageNet dataset.
+Figure 2 plots the results.
+We see from the figure that the Frobenius norm of of the $x^{3}$ archicture is much larger than softmax but scaling it by $\frac{1}{16}$ brings it down to softmax levels.
+Similarly, Figure 3 shows the Jacobian’s Frobenius norm, where scaling also brings the norms closer to softmax, ensuring more stable gradients.
+Further plots for other layers are in Section A.2.3.
+Table 1 presents the final Top-1% accuracy achieved by each activation function.
+Notably, $\frac{x^{3}}{16}$ delivers the best performance.
+In contrast, the unscaled $x^{3}$ activation yields significantly lower Top-1% accuracy, underscoring the importance of incorporating an appropriate scaling factor.
+Figure 2: Frobenius norm of the self-attention matrix with three different activations in layer 2, 7 and 12 of the ViT-Tiny architecture during training.
+Figure 3: Frobenius norm of the jacobian of the self-attention matrix with three different activations in layer 2, 7 and 12 of the ViT-Tiny architecture during training.
+Table 1: Comparison of Top-1% accuracy on Tiny-ImageNet between softmax and polynomial activations.
+softmax $\frac{x^{3}}{16}$ $x^{3}$
+Top-1% accuracy 50.26 50.5 45.3
 6 Experiments
-In this section, we validate the theory from section 4 on a variety of transformer tasks.
-We perform the empirical validation on two primary activations from section 4, namely a cubic polynomial activation $\phi(x)=x^{3}$ and a linear polynomial $\phi(x)=x$.
-The goal will be to show that by suitably scaling these activations using the theory in section 4, we can achieve competitive performance when compared to softmax.
-For the rest of this section we will simply denote these activations by $x^{3}$ and $x$.
-Figure 1: Training ViT-Tiny with the activation $\phi(x)=x^{3}$ with different sequence lengths and different scales.
-As the sequence length gets smaller, the log scale needed to obtain good accuracy decreases validating the theory from section 4.2.
+In this section, we evaluate a simple polynomial activation as an alternative to the standard softmax across a range of transformer applications commonly studied in the literature.
+The goal is to empirically challenge the conventional softmax properties, positivity, row-normalization, and sparsity—and examine whether these conditions are truly necessary for good transformer performance.
+Building on the theoretical foundations from Section 4, we focus on the cubic polynomial $x^{3}$ as a test case.
+Notably, $x^{3}$ introduces both positive and negative values, does not normalize rows to sum to 1, and generally produces dense attention matrices—violating all three traditional softmax conditions.
+We consider two scaling strategies for $x^{3}$:
+  • 1.
+    Fixed scale: Following Section 4, we scale $x^{3}$ by the inverse square root of the sequence length (which is fixed throughout training), as theory suggests this maintains optimization stability.
+  • 2.
+    Dynamic scale: Recognizing that the assumption of i.i.d.
+normal-distributed $Q$, $K$, $V$, used in Section 4, holds primarily at initialization, we also explore a learnable scale.
+This dynamic scale is initialized as above but optimized during training.
+Our experiments are not designed to achieve state-of-the-art results.
+Instead, they aim to question the softmax paradigm and show that alternative activations, even those violating softmax’s traditional properties, can still lead to effective transformer models.
+For each experiment we ran five trials and report the mean result.
 6.1 Image Classification
-5.1.1 ViT-Tiny on Tiny-Imagenet:
-In this section we test the theory from section 4 on the ViT-Tiny architecture (Steiner et al., 2021) trained from scratch on the Tiny-Imagenet dataset (Le & Yang, 2015).
-Our first experiment was to test how the Top-$1\%$ accuracy changes for a ViT-Tiny trained on Tiny-Imagenet as we change the sequence length of the input and the scale predicted in corollaries 4.3 and 4.5 when using the activation $x^{3}$.
-According to the theory developed in section 4, the Frobenius norm scales according to $\mathcal{O}(\sqrt{N})$ when we scale $X^{3}$ by $\frac{1}{\sqrt{N}}$.
-Thus as the sequence length decreases we should see the amount of scaling in a Log scale decrease.
-Figure 1 shows the results of this experiment.
-We considered four different input sequence lengths of sizes $256$, $64$, $16$ and $8$.
-We ran several ViT-Tiny architectures with a variety of scalings of the form $\mathcal{O}(\frac{1}{\sqrt{N}})$ where $N$ ranged below to above the sequence length.
-As can be seen from figure 1 as the sequence length got smaller the amount of scaling, shown in Log scale on the x-axis, needed for good accuracy got smaller verifying the theory in section 4.2.
-The second experiment compared activations $x^{3}$ and $x$, along with scaled versions $\frac{1}{16}x^{3}$ and $\frac{1}{16}x$, against softmax using the Tiny-ViT architecture on Tiny-Imagenet.
-With a sequence length of $256$ ($\sqrt{256}=16$), we decided to take $\frac{1}{16}$ as the scale of the polynomial activations.
-The experiment used a patch size of 4, 3 attention heads, and 12 layers as described in Steiner et al.
-(2021).
-Results in table 1 show $\frac{1}{8}x^{3}$ outperforming softmax, while the unscaled version performed poorly.
-Similarly, $\frac{1}{16}x$ performed competitively with a significant drop in performance without scaling.
-Figure 2 displays the Frobenius norm of the self-attention matrix during training for five activations in layers 2 and 12 of ViT-Tiny, averaged across all heads.
-Norms for $x^{3}$ and $x$ are higher than softmax, but scaling by $\frac{1}{16}$ reduces them to more stable levels, improving training stability.
-Similarly, figure 3 shows the Jacobian’s Frobenius norm, where scaling also brings the norms closer to softmax, ensuring more stable gradients.
-Further plots for other layers are in appendix A.2.2.
-softmax $\frac{x^{3}}{16}$ $x^{3}$ $\frac{x}{16}$ $x$
-Top-1% accuracy 50.26 50.5 45.3 47.9 31.78
-Table 1: Comparison of Top-1% accuracy on Tiny-Imagenet between softmax and polynomial activations.
-The cubic activation outperforms softmax when the right scale of $\frac{1}{8}$ is applied.
-Similarly, the linear activation is competitive only with an optimal scale.
-Figure 2: Frobenius norm of the self-attention matrix with five different activations in layer 2 and 12 of the ViT-Tiny architecture during training.
-Figure 3: Frobenius norm of the Jacobian of the self-attention matrix with five different activations in layer 2 and 12 of the ViT-Tiny architecture during training.
-5.1.2 Larger vision transformers on ImageNet-1k
-For this experiment we carried out an image classification task using a variety of different vision transformers from the literature on the ImageNet-1k dataset.
-We found that the scales $\frac{1}{14}$ worked best for both $x^{3}$ and $x$.
-We train all models on the ImageNet-1k dataset from scratch and report Top-1 accuracy on the validation set.
-We use PyTorch Paszke et al.
-(2019) and Timm Wightman (2019) libraries to train our models with similar setups to He et al.
-(2022) Liu et al.
-(2021).
-We examined our approach along with the following three transformer architectures to show its generalization:
-- ViT: Dosovitskiy et al.
-(2020) is the pioneering work that interprets an image as a sequence of patches and processes it by a standard Transformer encoder as used in NLP.
-This simple, yet scalable, strategy works surprisingly well when coupled with pre-training on large datasets.
-We use ViT-Small which has the following settings: patch size = 16, embedding dimensions = 384, number of heads = 6, and layers = 12.
-Also, We use ViT-Base which has the following settings: patch size = 16, embedding dimensions = 768, number of heads = 12, and layers = 12.
-- DeiT: Touvron et al.
-(2021) is a well-known transformer based on ViT.
-It is very similar to ViT except that it converged faster due to better training recipe since we do not use the distillation token imposed in DeiT.
-We use DeiT-Small which has the following settings: patch size = 16, embedding dimensions = 384, number of heads = 6, and layers = 12.
-Also, We use Deit-Base which has the following settings: patch size = 16, embedding dimensions = 768, number of heads = 12, and layers = 12.
-- Swin Transformer: Liu et al.
-(2021) produces a hierarchical feature representation and proposes the shifted window based self-attention which is shown to be effective and efficient on vision problems.
-We use Swin-Small with 96 channels and Swin-Base with 128 channels in the hidden layers of the first stage.
-The window size is set to $M$ = 7 by default, the query dimension of each head is $d$ = 32, and layer numbers are ${2,2,18,2}$ for all experiments.
-- XciT: Xiong et al.
-(2021) is a vision transformer architecture consisting of two different components compared to the standard ViT.
-Firstly, it has Local Patch Interaction in each block, which includes one depth-wise 3×3 convolution followed by Batch Normalization, GELU, and another depth-wise 3×3 convolution.
-Secondly, it uses Cross-Covariance attention, where the attention map is derived from the cross-covariance matrix computed over the key and query projections of the token features.
-We use XCiT-S12 with a patch size of 16 and XCiT-M24 with a patch size of 24.
+We conducted an image classification task using various vision transformer architectures from the literature on the ImageNet-1k dataset.
+For this task, the standard sequence length employed by the vision transformers on the ImageNet-1k dataset is $196$.
+We trained all models on the ImageNet-1k dataset from scratch and report Top-1 accuracy on the validation set.
+We examined our approach along with the following four transformer architectures to show its generalization, ViT-B [10], DeiT [25], Swin Transformer [17], XCiT [3].
+Each transformer was trained following the approach in each paper.
 The results are shown in Table 2.
-The activation $\frac{1}{14}x^{3}$ performed best on the ViT’s and the Swin transformers while softmax performed best on the DeiT architectures.
-Further ablations with different scales and activations can be fond in appendix A.2.1.
-Activation Models
-ViT-Base ViT-Small DeiT-Base DeiT-Small Swin-Base Swin-Small XciT-Medium XciT-Small
-softmax 79.6 80.2 78.9 79.6 83.0 83.3 81.1 81.2
-$\frac{x^{3}}{14}$ 79.6 80.5 77.4 78.3 83.2 83.4 81.2 82.1
-$x^{3}$ 77.8 78.6 76.1 76.3 79.7 79.9 78.1 78.3
-$\frac{x}{14}$ 76.9 77.8 79.6 79.8 79.4 79.5 79.2 79.3
-$x$ 73.2 73.9 77.7 77.9 77.8 77.9 76.4 76.6
-Table 2: Comparsions of pre-training models with different activation functions on ImageNet-1k datasets. We report the classification top-1 accuracy (%).
-Figure 4 plots the Frobenius norm of the self-attention matrix with the five different activations during training in layers 2 and 12, averaged over all heads within each layer, of the ViT-Small architecture.
-By scaling the activations $x^{3}$ and $x$ by $\frac{1}{14}$ we were able to control the scale of the Frobenius norm of the self-attention matrix and obtain scales comparable to softmax’s scale.
-Simiarly, figure 5 plots the Frobenius norm of the Jacobian of the self-attention matrix during training for layers 2 and 12, averaged over all heads.
-By scaling the activations $x^{3}$ and $x$ by $\frac{1}{14}$ we were able to control the scale of the Jacobian norm and obtain scales comparable to softmax’s scale during training.
-Plots for other layers of the architecture during training are given in appendix A.2.2.
-Figure 4: Frobenius norm of the self-attention matrix with five different activations in layer 2 and 12 of the ViT-Tiny architecture during training.
-Figure 5: Frobenius norm of the Jacobian of the self-attention matrix with five different activations in layer 2 and 12 of the ViT-Tiny architecture during training.
-5.1.3 Visualizing self-attention with ViT-Base
-We plotted heat maps of self-attention matrices using the $\frac{1}{14}x^{3}$ and softmax activations across two layers and heads after convergence, averaging over a fixed training batch of size 128.
-Figure 6 shows layer 2, head 8, highlighting the differences in attention patterns, with $\frac{1}{14}x^{3}$ containing both positive and negative values.
-Similarly, Figure 7 for layer 12, head 6 shows distinct patterns for each activation.
-Overall, the ViT-Base architecture with $\frac{1}{14}x^{3}$ exhibits notably different self-attention patterns compared to softmax.
-We visualized how the self-attention matrix targets different regions of an image.
-Using an image from the ImageNet-1k validation set, we extracted the class token, reshaped it into a $(14,14)$ grid representing 196 patches, and then mapped it back to the original image size with nearest neighbor interpolation.
-Figure 8 shows the input image, while figure 9 illustrates the self-attention matrices for different activations in layer 12, head 6 of the ViT-Base architecture after convergence, highlighting their distinct focus areas.
-Figure 6: Heat maps of the self-attention matrix in layer 2, head 8, of a ViT base architecture, comparing $\frac{1}{14}x^{3}$ (left) and softmax (right) activations after training.
+As can be seen from the table the $x^{3}$ activation with a dynamic scale performed the best and the one with a fixed scale performed comparable to softmax.
+On the other hand $x^{3}$ (with no scale) underperformed on all ViTs showing how important the theory on scaling as developed in Section 4 is.
+For a discussion on other polynomials and linear attention see Section A.2.4.
+Table 2: Comparisons of pre-training models with different activation functions on ImageNet-1k. We report top-1 classification accuracy (%).
+Models
+ViT-B DeiT-B Swin-B XciT-M
+softmax 80.3 81.5 83.5 82.7
+$x^{3}$ + fixed scale 80.2 81.4 83.6 82.8
+$x^{3}$ + dynamic scale 80.3 81.6 83.6 82.9
+$x^{3}$ 78.1 78.5 79.9 79.5
+Visualizing attention heads:
+Softmax attention traditionally satisfies three key properties: positivity, row normalization, and sparsity.
+In contrast, the polynomial activation $x^{3}$ (with or without positive scaling) takes positive values for $x>0$ and negative values for $x<0$, thus violating these softmax constraints.
+To better understand how this affects attention patterns, we analyzed the self-attention matrices of ViT-B models trained with the $x^{3}$ + dynamic scale activation and compared them to those using softmax.
+We visualized heatmaps of the attention matrices after convergence, focusing on two representative layers and heads, and averaged over a fixed batch of 128 samples.
+Figure 4 shows results for layer 2, head 8, where the $x^{3}$ + dynamic scale activation produces attention scores with both positive and negative values, unlike softmax.
+Similarly, Figure 5 illustrates distinct patterns for layer 12, head 6, highlighting how the two activations differ in learned attention distributions.
+Interpretability.
+Even with these differences, the $x^{3}$ + dynamic scale activation still learns meaningful attention patterns.
+For instance, in Figure 4 (left), we observe 14 bands about the diagonal, indicating that the head has learned to attend to patches in the same image row as the query patch, sufficient for effective image classification.
+This suggests that the sign of attention values (positive or negative) is not inherently critical for attention allocation.
+Further, Figure 5 (left) reveals vertical lines, showing attention that depends solely on key positions, independent of the query position.
+Smaller key indices receive higher attention weights, focusing model capacity where it matters most for classification tasks.
+These findings demonstrate that, despite deviating from softmax properties, the $x^{3}$ + dynamic scale activation enables the model to discover effective attention patterns.
+We noticed similar attention patterns for the $x^{3}$ + fixed scale activation.
+Figure 4: Heat maps of the self-attention matrix in layer 2, head 8, of a ViT base architecture, comparing $x^{3}$ + dynamic (left) and softmax (right) activations after training.
 The stark difference in self-attention patterns between the two activations is evident, showing distinct distributions across input tokens.
-Figure 7: Heat maps of the self-attention matrix in layer 12, head 6, of a ViT base architecture, comparing $\frac{1}{14}x^{3}$ (left) and softmax (right) activations after training.
+Figure 5: Heat maps of the self-attention matrix in layer 12, head 6, of a ViT base architecture, comparing $x^{3}$ + dynamic scale (left) and softmax (right) activations after training.
 The contrast in self-attention patterns between the two activations is clearly visible.
-Figure 8: Fish image from validation set of ImageNet-1k broken up into $196=14\times 14$ patches.
-Figure 9: Comparing how $\frac{1}{14}x^{3}$ and softmax self-attention matrices focus on different parts of the fish image from figure 8 in layer 12, head 6, after the model has converged.
 6.2 Object Detection and Instance Segmentation
-In this section, in order to examine the transfer learning ability of our models, we demonstrate our approach to object detection and segmentation tasks by fine-tuning our ImageNet-pretrained XCiT model on them.
-Our experiments are conducted on COCO 2017Lin et al.
-(2014), which has 118K training images and 5K validation images with 80 categories.
-We integrate the XCiT architecture as the backbone in the Mask R-CNN (He et al., 2017) detector with a Feature Pyramid Network (FPN).
+In this section, in order to examine the transfer learning ability of our models, we demonstrate our approach on object detection and segmentation tasks by fine-tuning a ImageNet-pretrained XCiT model on them.
+Our experiments are conducted on COCO 2017 [16], which has 118K training images and 5K validation images with 80 categories.
+We integrate the XCiT architecture as the backbone in the Mask R-CNN [13] detector with a Feature Pyramid Network (FPN).
 Due to XCiT’s inherently columnar design, we adapt it for FPN compatibility by extracting features from various layers for XCiT-S12.
-These features have a consistent stride of either 16.
-The feature resolutions are then adjusted to strides of [4, 8, 16, 32] This downsampling is accomplished through max pooling, while upsampling is achieved using a single transposed convolution layer.
-The model is trained for 36 epochs using the AdamW optimizer with learning rate of $10^{-4}$, 0.05 weight decay and 16 batch size.
-In table 3,we condunct experiments on XCiT-S12 models using 16×16 patches with the activations $\frac{1}{14}x^{3}$, $\frac{1}{14}x$ and softmax.
-We found we couldn’t train with the activations $x^{3}$ and $x$ on this task well so only report the others.
-Activation $AP^{b}$ $AP^{b}_{50}$ $AP^{b}_{75}$ $AP^{m}$ $AP^{m}_{50}$ $AP^{m}_{75}$
-softmax 44.9 66.1 48.9 40.1 63.1 42.8
-$\frac{x^{3}}{14}$ 44.8 66.3 49 40.1 63.1 42.9
-$\frac{x}{14}$ 44.8 66.2 49 40.1 63.1 42.8
+These features have a consistent stride of 16.
+The feature resolutions are then adjusted to strides of [4, 8, 16, 32].
+This downsampling is accomplished through max pooling, while upsampling is achieved using a single transposed convolution layer.
+We conducted experiments on XCiT-S12 models using 16×16 patches with the activations $x^{3}$ + fixed scale, $x^{3}$ + dynamic scale, with scale initialized to $\frac{1}{14}$, and softmax.
+We found when we trained with $x^{3}$ the model would not converge thereby suggesting that a scaling is extremely important as predicted by the theory in Section 4.2.
+The training hyperparameters are given in Section A.2.2.
+Table 3 shows the results of the experiment.
+From the table we see that the activation $x^{3}$ + dynamic scale obtains better performance to softmax with $x^{3}$ + fixed scale obtaining comparable performance.
 Table 3: COCO object detection and instance segmentation performance on the mini-val set.
 All backbones are pretrained on ImageNet-1k, and use Mask R-CNN model.
-$AP^{b}$: Average Precision for bounding box predictions, $AP^{b}_{50/75}$: Average Precision at an Intersection over Union (IoU) threshold of 0.50/0.75 for bounding box predictions, $AP^{m}$: Average Precision for mask predictions, $AP^{m}_{50/75}$: Average Precision at an Intersection over Union (IoU) threshold of 0.50/0.75 for mask predictions
-5.3 Natural language processing(NLP)
-To assess the effectiveness of our approach on NLP tasks, we trained models on five benchmarks from the Long Range Arena (LRA) suite Tay et al.
-(2020): ListOps, Text Classification, Retrieval, Image Classification, and Pathfinder.
-We evaluated the activations $\frac{x^{3}}{14}$ and $\frac{x}{14}$ against softmax, finding that $x^{3}$ and $x$ did not train effectively on their own, so only results for these scaled activations and softmax are presented.
-Our implementation followed the guidelines from Xiong et al.
-(2021).
-The results are summarized in table 4.
-Activation ListOps Text Retrieval Image Pathfinder
+$AP^{b}$: Average Precision for bounding box predictions, $AP^{b}_{50/75}$: Average Precision at an IoU threshold of 0.50/0.75 for bounding box predictions, $AP^{m}$: Average Precision for mask predictions, $AP^{m}_{50/75}$: Average Precision at an IoU threshold of 0.50/0.75 for mask predictions.
+$AP^{b}$ $AP^{b}_{50}$ $AP^{b}_{75}$ $AP^{m}$ $AP^{m}_{50}$ $AP^{m}_{75}$
+softmax 44.9 66.1 48.9 40.1 63.1 42.8
+$x^{3}$ + fixed scale 44.8 66.3 49.1 40.2 63.1 43.0
+$x^{3}$ + dynamic scale 45.1 66.5 49.4 40.4 63.2 43.1
+6.3 Nyströmformer on LRA benchmark
+Nyströmformer is a transformer designed to handle long-range dependencies more efficiently by approximating the self-attention mechanism with the Nyström method [30].
+To assess the effectiveness of our approach, we trained models on five benchmarks from the Long Range Arena (LRA) suite [23]: ListOps, Text Classification, Retrieval, Image Classification, and Pathfinder.
+On each dataset we compared a Nyströmformer with softmax, $x^{3}$ + dynamic scale, and $x^{3}$.
+For the $x^{3}$ + dynamic scale activation on each dataset we used a different scale as initialization based on the sequence length used in the original Nyströmformer model.
+For ListOps the original model used a sequence length of $512$ and so the scale we used as initialization was $1/\sqrt{512}$, on Text Classification the input sequence used is $1024$ and so the scale we used at initialization was $1/\sqrt{1024}$, on Retrieval the input sequence used is $2048$ and the initialized scale we used was $1/\sqrt{2048}$, on Image Classification the input sequence is $4096$ and the initialized scale we used was $1/\sqrt{4096}$, and on Pathfinder the input sequence used is $8192$ and the initialized scale we used was $1/\sqrt{8192}$.
+We followed the training regime of the original paper [30].
+The results are shown in Table 4.
+As can be seen from the table the $x^{3}$ + dynamic scale activation results in better performance compared to softmax with $x^{3}$ + fixed scale performing comparably.
+However $x^{3}$ (with no scale) underperforms on all tasks.
+Table 4: Comparisons of Nyströmformer models with different activation functions on the LRA benchmark. We report the accuracy (%).
+ListOps Text Retrieval Image Pathfinder
 softmax 37.1 63.8 79.8 39.9 72.9
-$\frac{x^{3}}{14}$ 37.5 63.3 80.9 37.2 68.7
-$\frac{x}{14}$ 34.3 62.9 81.5 39.0 69.1
-Table 4: Comparsions of transformer models with different activation functions on NLP tasks. We report the accuracy (%) on LRA benchmarks.
+$x^{3}$ + fixed scale 37.2 63.7 80.0 39.9 72.9
+$x^{3}$ + dynamic scale 37.5 63.9 81.1 40.1 73.1
+$x^{3}$ 32.3 62.0 78.5 38.1 67.9
+7 Physics Informed Transformers
+Recently, dynamical data has become essential for modeling time-dependent systems.
+[31] introduced transformer architectures for solving partial differential equations (PDEs), demonstrating superior performance across various PDE tasks.
+These models effectively capture interactions between spatial and temporal components inherent in dynamical systems.
+In Section A.2.5, we extend our polynomial attention framework to Physics-Informed Transformers, showing that polynomial activations can outperform softmax in this setting.
 8 Limitations
-While our work introduces novel activations that challenge the conventional softmax approach, there are some limitations to address.
-Our theoretical framework is primarily designed for dot-product self-attention and may not immediately extend to other attention mechanisms, although our empirical results showed competitive performance against softmax across different architectures.
-Additionally, we observed that while our activations performed well on vision tasks, their performance was less consistent on NLP tasks, suggesting that a more refined theoretical approach may be needed for these applications.
+While our work introduces polynomial activations that challenge the conventional understanding of softmax, there are several limitations to address.
+First, our theoretical framework is specifically developed for dot-product self-attention and may not directly generalize to other forms of attention mechanisms, such as additive attention or kernelized approximations.
+Extending our analysis to these variants could reveal additional nuances.
+Second, while our empirical evaluations span multiple architectures and tasks, they are limited to models with up to 100 million parameters due to resource constraints.
+The scalability of polynomial activations in large-scale transformers, especially those with billions of parameters, should be considered in the future.
 9 Conclusion
-This work challenges the traditional view that transformer activations for attention must produce sparse probability distributions.
+This work questioned whether transformer activations for attention must produce sparse probability distributions.
 We introduced a theoretical framework analyzing the Frobenius norm of the self-attention matrix, which suggests key scaling properties for activations in attention mechanisms.
 We proved that specific polynomial activations, which behave very differently from softmax, satisfy these properties.
-Through extensive experiments across vision and NLP tasks, we demonstrated that these alternative activations not only compete with but sometimes outperform softmax, offering a fresh perspective on attention mechanisms in transformers.
+Through extensive experiments across a variety of transformer applications, we demonstrated that these alternative activations not only compete with but can outperform softmax, offering a new perspective on attention mechanisms in transformers.
 Appendix A Appendix
@@
 We will start by proving the first inequality in theorem 4.1. Given a matrix $A=(a_{ij})\in\mathbb{R}^{N\times N}$ we have that
-$\mathbf{softmax}\bigg{(}\begin{bmatrix}a_{11}&\cdots&a_{1n}\\ \vdots&\vdots&\vdots\\ a_{n1}&\cdots&a_{nn}\end{bmatrix}\bigg{)}=\begin{bmatrix}\frac{e^{a_{11}}}{\sum_{j=1}^{n}e^{a_{1j}}}&\cdots&\frac{e^{a_{1n}}}{\sum_{j=1}^{n}e^{a_{1j}}}\\ \vdots&\vdots&\vdots\\ \frac{e^{a_{n1}}}{\sum_{j=1}^{n}e^{a_{nj}}}&\cdots&\frac{e^{a_{nn}}}{\sum_{j=1}^{n}e^{a_{nj}}}.\end{bmatrix}$ (A.1)
+$\mathbf{softmax}\bigg{(}\begin{bmatrix}a_{11}&\cdots&a_{1n}\\ \vdots&\vdots&\vdots\\ a_{n1}&\cdots&a_{nn}\end{bmatrix}\bigg{)}=\begin{bmatrix}\frac{e^{a_{11}}}{\sum_{j=1}^{n}e^{a_{1j}}}&\cdots&\frac{e^{a_{1n}}}{\sum_{j=1}^{n}e^{a_{1j}}}\\ \vdots&\vdots&\vdots\\ \frac{e^{a_{n1}}}{\sum_{j=1}^{n}e^{a_{nj}}}&\cdots&\frac{e^{a_{nn}}}{\sum_{j=1}^{n}e^{a_{nj}}}.\end{bmatrix}$ (12)
 By definition of the Frobenius norm we then see that
-$||\mathbf{softmax}(A)||_{F}^{2}$ $=\bigg{(}\frac{1}{\sum_{j=1}^{n}e^{a_{1j}}}\bigg{)}^{2}(e^{2a_{11}}+\cdots e^{2a_{11}})+\cdots+\bigg{(}\frac{1}{\sum_{j=1}^{n}e^{a_{Nj}}}\bigg{)}^{2}(e^{2a_{N1}}+\cdots e^{2a_{NN}})$ (A.2)
-$\leq\bigg{[}\bigg{(}\frac{1}{\sum_{j=1}^{n}e^{a_{1j}}}\bigg{)}(e^{a_{11}}+\cdots e^{a_{11}})\bigg{]}^{2}+\cdots+\bigg{[}\bigg{(}\frac{1}{\sum_{j=1}^{n}e^{a_{Nj}}}\bigg{)}(e^{a_{N1}}+\cdots e^{a_{NN}})\bigg{]}^{2}$ (A.3)
-$=1+\cdots+1$ (A.4)
-$=N$ (A.5)
+$||\mathbf{softmax}(A)||_{F}^{2}$ $=\bigg{(}\frac{1}{\sum_{j=1}^{n}e^{a_{1j}}}\bigg{)}^{2}(e^{2a_{11}}+\cdots e^{2a_{11}})+\cdots$ (13)
+$\hskip 28.45274pt+\bigg{(}\frac{1}{\sum_{j=1}^{n}e^{a_{Nj}}}\bigg{)}^{2}(e^{2a_{N1}}+\cdots e^{2a_{NN}})$ (14)
+$\leq\bigg{[}\bigg{(}\frac{1}{\sum_{j=1}^{n}e^{a_{1j}}}\bigg{)}(e^{a_{11}}+\cdots e^{a_{11}})\bigg{]}^{2}+\cdots$ (15)
+$\hskip 28.45274pt+\bigg{[}\bigg{(}\frac{1}{\sum_{j=1}^{n}e^{a_{Nj}}}\bigg{)}(e^{a_{N1}}+\cdots e^{a_{NN}})\bigg{]}^{2}$ (16)
+$=1+\cdots+1$ (17)
+$=N$ (18)
 where the second inequality uses the fact that for non-negative numbers $a$ and $b$ we always have that $a^{2}+b^{2}\leq(a+b)^{2}$.
 It then immediately follows that $||\mathbf{softmax}(A)||_{F}\leq\sqrt{N}$ and this proves the first inequality in the statement of theorem 4.1.
-We move on to prove the second inequality in the statement of theorem 4.1. For this, let us write each entry of the matrix on the right of equation A.1 as follows:
-$F_{kl}=\frac{e^{a_{kl}}}{\sum_{j=1}^{N}e^{a_{kj}}}.$ (A.6)
+We move on to prove the second inequality in the statement of theorem 4.1. For this, let us write each entry of the matrix on the right of (12) as follows:
+$F_{kl}=\frac{e^{a_{kl}}}{\sum_{j=1}^{N}e^{a_{kj}}}.$ (19)
 By applying the chain rule we then have the following derivative formulas
-$\frac{\partial}{\partial x_{ij}}F_{ij}$ $=F_{ij}-F_{ij}^{2}$ (A.7)
-$\frac{\partial}{\partial x_{ik}}F_{ij}$ $=-F_{ij}F_{ik}\text{ for any }k\neq j$ (A.8)
-$\frac{\partial}{\partial x_{kl}}F_{ij}$ $=0\text{ for any }k\neq i\text{ and }l\neq j.$ (A.9)
+$\frac{\partial}{\partial x_{ij}}F_{ij}$ $=F_{ij}-F_{ij}^{2}$ (20)
+$\frac{\partial}{\partial x_{ik}}F_{ij}$ $=-F_{ij}F_{ik}\text{ for any }k\neq j$ (21)
+$\frac{\partial}{\partial x_{kl}}F_{ij}$ $=0\text{ for any }k\neq i\text{ and }l\neq j.$ (22)
 We can then express the gradient as
-$\nabla\mathbf{softmax}(A)=\begin{bmatrix}\nabla F_{11}\\ \vdots\\ \nabla F_{1N}\\ \nabla F_{21}\\ \vdots\\ \vdots\\ \nabla F_{NN}\end{bmatrix}$ (A.10)
+$\nabla\mathbf{softmax}(A)=\begin{bmatrix}\nabla F_{11}\\ \vdots\\ \nabla F_{1N}\\ \nabla F_{21}\\ \vdots\\ \vdots\\ \nabla F_{NN}\end{bmatrix}$ (23)
 where
-$\nabla F_{ij}=\begin{bmatrix}-F_{ij}F_{i1}^{2}&-F_{ij}F_{i2}&\cdots&F_{ij}-F_{ij}^{2}&\cdots&F_{ij}F_{iN}.\end{bmatrix}$ (A.11)
+$\nabla F_{ij}=\begin{bmatrix}-F_{ij}F_{i1}^{2}&-F_{ij}F_{i2}&\cdots&F_{ij}-F_{ij}^{2}&\cdots&F_{ij}F_{iN}.\end{bmatrix}$ (24)
 From these computations we see that
-$||\nabla\mathbf{softmax}(A)||_{F}^{2}=||\nabla F_{11}||_{F}^{2}+\cdots+||\nabla F_{NN}||_{F}^{2}.$ (A.12)
+$||\nabla\mathbf{softmax}(A)||_{F}^{2}=||\nabla F_{11}||_{F}^{2}+\cdots+||\nabla F_{NN}||_{F}^{2}.$ (25)
 We will proceed by bounding each collection $||\nabla F_{i1}||_{F}^{2}+\cdots+||\nabla F_{1N}||_{F}^{2}$ separately then add up all the bounds. We have
-$||\nabla F_{i1}||_{F}^{2}+\cdots+||\nabla F_{1N}||_{F}^{2}$ $=|F_{i1}-F_{i1}^{2}|^{2}+|F_{i1}F_{i2}|^{2}+\cdots+|F_{i1}F_{iN}|^{2}$ (A.13)
-$\hskip 14.22636pt+|F_{i2}F_{i1}|^{2}+|F_{i2}-F_{i2}^{2}|^{2}+\cdots+|F_{i2}F_{iN}|^{2}$ (A.14)
+$||\nabla F_{i1}||_{F}^{2}+\cdots+||\nabla F_{1N}||_{F}^{2}$ $=|F_{i1}-F_{i1}^{2}|^{2}+|F_{i1}F_{i2}|^{2}+\cdots+|F_{i1}F_{iN}|^{2}$ (26)
+$\hskip 14.22636pt+|F_{i2}F_{i1}|^{2}+|F_{i2}-F_{i2}^{2}|^{2}+\cdots+|F_{i2}F_{iN}|^{2}$ (27)
 […]
-$\hskip 14.22636pt+|F_{iN}F_{i1}|^{2}+|F_{iN}F_{i2}|^{2}+\cdots+|F_{iN}-F_{iN}^{2}|^{2}$ (A.16)
-$\leq(F_{i1})^{2}(|1-F_{i1}|+|F_{i2}|+\cdots+|F_{iN}|)^{2}$ (A.17)
-$\hskip 14.22636pt(F_{i2})^{2}(|F_{i1}|+|1-F_{i2}|+\cdots+|F_{iN}|)^{2}$ (A.18)
+$\hskip 14.22636pt+|F_{iN}F_{i1}|^{2}+|F_{iN}F_{i2}|^{2}+\cdots+|F_{iN}-F_{iN}^{2}|^{2}$ (29)
+$\leq(F_{i1})^{2}(|1-F_{i1}|+|F_{i2}|+\cdots+|F_{iN}|)^{2}$ (30)
+$\hskip 14.22636pt(F_{i2})^{2}(|F_{i1}|+|1-F_{i2}|+\cdots+|F_{iN}|)^{2}$ (31)
 […]
-$\hskip 14.22636pt+(F_{iN})^{2}(|F_{i1}|+|F_{i2}|+\cdots+|1-F_{iN}|)^{2}.$ (A.20)
+$\hskip 14.22636pt+(F_{iN})^{2}(|F_{i1}|+|F_{i2}|+\cdots+|1-F_{iN}|)^{2}.$ (33)
 We then observe that since $F_{i1}+\cdots+F_{iN}=1$ we have that $1-F_{ij}=2(F_{i1}+\cdots+\widehat{F_{ij}}+\cdots+F_{iN})$ where $\widehat{F_{ij}}$ means we don’t include $F_{ij}$ in the sum.
 This means we get the bound
-$||\nabla F_{i1}||_{F}^{2}+\cdots+||\nabla F_{1N}||_{F}^{2}$ $\leq 4F_{i1}^{2}(\widehat{F_{i1}}+F_{i2}+\cdots+F_{iN})$ (A.21)
+$||\nabla F_{i1}||_{F}^{2}+\cdots+||\nabla F_{1N}||_{F}^{2}$ $\leq 4F_{i1}^{2}(\widehat{F_{i1}}+F_{i2}+\cdots+F_{iN})$ (34)
 […]
-$\hskip 14.22636pt+4F_{iN}^{2}(F_{i1}+F_{i2}+\cdots+\widehat{F_{iN}})$ (A.23)
-$\leq 4(F_{i1}^{2}+\cdots F_{iN}^{2})$ (A.24)
-$=4.$ (A.25)
+$\hskip 14.22636pt+4F_{iN}^{2}(F_{i1}+F_{i2}+\cdots+\widehat{F_{iN}})$ (36)
+$\leq 4(F_{i1}^{2}+\cdots F_{iN}^{2})$ (37)
+$=4.$ (38)
 Putting all the bounds together for each of the terms $N$ terms $||\nabla F_{i1}||_{F}^{2}+\cdots+||\nabla F_{1N}||_{F}^{2}$ we get
-$||\nabla\mathbf{softmax}(A)||_{F}^{2}\leq 4N$ (A.26)
+$||\nabla\mathbf{softmax}(A)||_{F}^{2}\leq 4N$ (39)
 and this implies
-$||\nabla\mathbf{softmax}(A)||_{F}\leq 2\sqrt{N}.$ (A.27)
+$||\nabla\mathbf{softmax}(A)||_{F}\leq 2\sqrt{N}.$ (40)
 This finishes the proof of theorem 4.1.
@@
 For the $p=1$ case we can then compute
-$\mathop{\mathbb{E}}(c_{11}^{2})$ $=\mathop{\mathbb{E}}((\sum_{i=1}^{D}a_{1i}b_{i1})^{2})$ (A.28)
+$\mathop{\mathbb{E}}(c_{11}^{2})$ $=\mathop{\mathbb{E}}((\sum_{i=1}^{D}a_{1i}b_{i1})^{2})$ (41)
 $=\mathop{\mathbb{E}}(\sum_{i=1}^{D}a_{1i}^{2}b_{i1}^{2}+\sum_{i=1}^{D}\sum_{j=1,j\neq i}^{D}a_{1i}b_{i1}a_{1j}b_{j1})$
@@
 We will compute the expectation of the first entry $c_{1}1\in\mathbf{C}$:
-$\mathop{\mathbb{E}}(c_{11}^{2})$ $=\mathop{\mathbb{E}}((\sum_{i=1}^{D}a_{1i}b_{i1})^{2p})$ (A.29)
+$\mathop{\mathbb{E}}(c_{11}^{2})$ $=\mathop{\mathbb{E}}((\sum_{i=1}^{D}a_{1i}b_{i1})^{2p})$ (42)
 $=\mathop{\mathbb{E}}(\sum_{i=1}^{D}a_{1i}^{2p}b_{i1}^{2p}+\sum_{i=1}^{D}\sum_{j=1,j\neq i}^{D}a_{1i}^{2p-2}b_{i1}^{2p-2}a_{1j}^{2}b_{j1}^{2}+\cdots).$
@@
 The count is similar to choosing $p$ items from $D$,
-$\mathop{\mathbb{E}}(c_{11}^{2})$ $\approx\mathop{\mathbb{E}}(\sum_{\{i_{1},\dots,i_{p}\}\in\{1,\dots,D\}}\prod_{k=1}^{p}a_{1,i_{k}}^{2}b_{i_{k},1}^{2})$ (A.30)
+$\mathop{\mathbb{E}}(c_{11}^{2})$ $\approx\mathop{\mathbb{E}}(\sum_{\{i_{1},\dots,i_{p}\}\in\{1,\dots,D\}}\prod_{k=1}^{p}a_{1,i_{k}}^{2}b_{i_{k},1}^{2})$ (43)
 $=\left(\begin{array}[]{c}D\\ p\end{array}\right)\frac{2p!}{2^{p}}\sigma_{1}^{2p}\sigma_{2}^{2p}$
@@
 We will do the $p=1$ case first. We proceed similar to the proof of Theorem 4.2.
-$\mathop{\mathbb{E}}(\|\frac{\partial XQK^{T}X^{T}}{\partial Q}\|^{2}_{F})$ $=\sum_{i=1}^{N}\sum_{j=1}^{N}\mathop{\mathbb{E}}(|\frac{\partial x_{i}^{T}QK^{T}x_{j}}{\partial Q}|^{2}_{F})$ (A.31)
-$=\sum_{i=1}^{N}\sum_{j=1}^{N}\mathop{\mathbb{E}}(\|x_{i}x_{j}^{T}K\|^{2}_{F})$ (A.32)
-$=\sum_{i=1}^{N}\sum_{j=1}^{N}\mathop{\mathbb{E}}(\sum_{k=1}^{D}\sum_{l=1}^{d}(x_{ik}\sum_{m=1}^{D}x_{jm}k_{ml})^{2})$ (A.33)
-$=\sum_{i=1}^{N}\sum_{j=1}^{N}\mathop{\mathbb{E}}(\sum_{k=1}^{D}\sum_{l=1}^{d}x_{ik}^{2}(\sum_{m=1}^{D}x_{jm}k_{ml})^{2})$ (A.34)
-$=\sum_{i=1}^{N}\sum_{j=1}^{N}\mathop{\mathbb{E}}(\sum_{k=1}^{D}\sum_{l=1}^{d}x_{ik}^{2}(\sum_{m=1}^{D}x_{jm}^{2}k_{ml}^{2}+\sum_{m=1}^{D}\sum_{n=1,n\neq m}^{D}x_{jm}k_{ml}x_{jn}k_{nl}))$ (A.35)
-$=\sum_{i=1}^{N}\sum_{j=1}^{N}\sum_{k=1}^{D}\sum_{l=1}^{d}(\sum_{m=1}^{D}\mathop{\mathbb{E}}(x_{ik}^{2}x_{jm}^{2}k_{ml}^{2})+\sum_{m=1}^{D}\sum_{n=1,n\neq m}^{D}\mathop{\mathbb{E}}(x_{ik}^{2}x_{jm}k_{ml}x_{jn}k_{nl}))$ (A.36)
-$=\sum_{i=1}^{N}\sum_{j=1}^{N}\sum_{k=1}^{D}\sum_{l=1}^{d}(\sum_{m=1}^{D}\mathop{\mathbb{E}}(x_{ik}^{2}x_{jm}^{2}k_{ml}^{2})+0)$ (A.37)
-$=\sum_{i=1}^{N}\sum_{j=1}^{N}\sum_{k=1}^{D}\sum_{l=1}^{d}\sum_{m=1}^{D}\mathop{\mathbb{E}}(x_{ik}^{2}x_{jm}^{2}k_{ml}^{2})$ (A.38)
-$=\sum_{i=1}^{N}\sum_{k=1}^{D}\sum_{l=1}^{d}\mathop{\mathbb{E}}(x_{ik}^{2}x_{ik}^{2}k_{kl}^{2})+\sum_{i=1}^{N}\sum_{j=1,j\neq i}^{N}\sum_{k=1}^{D}\sum_{l=1}^{d}\sum_{m=1,m\neq k}^{D}\mathop{\mathbb{E}}(x_{ik}^{2}x_{jm}^{2}k_{ml}^{2})$ (A.39)
-$=NDd3\sigma_{x}^{4}\sigma_{w}^{2}+N(N-1)D(D-1)d\sigma_{x}^{4}\sigma_{w}^{2}$ (A.40)
-$\approx N^{2}D^{2}d\sigma_{x}^{4}\sigma_{w}^{2}.$ (A.41)
+$\mathop{\mathbb{E}}(\|\frac{\partial XQK^{T}X^{T}}{\partial Q}\|^{2}_{F})$ $=\sum_{i=1}^{N}\sum_{j=1}^{N}\mathop{\mathbb{E}}(|\frac{\partial x_{i}^{T}QK^{T}x_{j}}{\partial Q}|^{2}_{F})$ (44)
+$=\sum_{i=1}^{N}\sum_{j=1}^{N}\mathop{\mathbb{E}}(\|x_{i}x_{j}^{T}K\|^{2}_{F})$ (45)
+$=\sum_{i=1}^{N}\sum_{j=1}^{N}\mathop{\mathbb{E}}(\sum_{k=1}^{D}\sum_{l=1}^{d}(x_{ik}\sum_{m=1}^{D}x_{jm}k_{ml})^{2})$ (46)
+$=\sum_{i=1}^{N}\sum_{j=1}^{N}\mathop{\mathbb{E}}(\sum_{k=1}^{D}\sum_{l=1}^{d}x_{ik}^{2}(\sum_{m=1}^{D}x_{jm}k_{ml})^{2})$ (47)
+$=\sum_{i=1}^{N}\sum_{j=1}^{N}\mathop{\mathbb{E}}(\sum_{k=1}^{D}\sum_{l=1}^{d}x_{ik}^{2}(\sum_{m=1}^{D}x_{jm}^{2}k_{ml}^{2}+\sum_{m=1}^{D}\sum_{n=1,n\neq m}^{D}x_{jm}k_{ml}x_{jn}k_{nl}))$ (48)
+$=\sum_{i=1}^{N}\sum_{j=1}^{N}\sum_{k=1}^{D}\sum_{l=1}^{d}(\sum_{m=1}^{D}\mathop{\mathbb{E}}(x_{ik}^{2}x_{jm}^{2}k_{ml}^{2})$ (49)
+$\hskip 28.45274pt+\sum_{m=1}^{D}\sum_{n=1,n\neq m}^{D}\mathop{\mathbb{E}}(x_{ik}^{2}x_{jm}k_{ml}x_{jn}k_{nl}))$ (50)
+$=\sum_{i=1}^{N}\sum_{j=1}^{N}\sum_{k=1}^{D}\sum_{l=1}^{d}(\sum_{m=1}^{D}\mathop{\mathbb{E}}(x_{ik}^{2}x_{jm}^{2}k_{ml}^{2})+0)$ (51)
+$=\sum_{i=1}^{N}\sum_{j=1}^{N}\sum_{k=1}^{D}\sum_{l=1}^{d}\sum_{m=1}^{D}\mathop{\mathbb{E}}(x_{ik}^{2}x_{jm}^{2}k_{ml}^{2})$ (52)
+$=\sum_{i=1}^{N}\sum_{k=1}^{D}\sum_{l=1}^{d}\mathop{\mathbb{E}}(x_{ik}^{2}x_{ik}^{2}k_{kl}^{2})$ (53)
+$\hskip 28.45274pt+\sum_{i=1}^{N}\sum_{j=1,j\neq i}^{N}\sum_{k=1}^{D}\sum_{l=1}^{d}\sum_{m=1,m\neq k}^{D}\mathop{\mathbb{E}}(x_{ik}^{2}x_{jm}^{2}k_{ml}^{2})$ (54)
+$=NDd3\sigma_{x}^{4}\sigma_{w}^{2}+N(N-1)D(D-1)d\sigma_{x}^{4}\sigma_{w}^{2}$ (55)
+$\approx N^{2}D^{2}d\sigma_{x}^{4}\sigma_{w}^{2}.$ (56)
 When $p>1$ we can proceed in a similar way.
-$\mathop{\mathbb{E}}(\|\frac{\partial(XQK^{T}X^{T})^{p}}{\partial Q}\|^{2}_{F})$ $=\sum_{i=1}^{N}\sum_{j=1}^{N}\mathop{\mathbb{E}}(\|\frac{(\partial x_{i}^{T}QK^{T}x_{j})^{p}}{\partial Q}\|^{2}_{F})$ (A.42)
-$=\sum_{i=1}^{N}\sum_{j=1}^{N}\mathop{\mathbb{E}}(\|p(x_{i}^{T}QK^{T}x_{j})^{p-1}\frac{\partial x_{i}^{T}QK^{T}x_{j}}{\partial Q}\|^{2}_{F})$ (A.43)
-$=\sum_{i=1}^{N}\sum_{j=1}^{N}\mathop{\mathbb{E}}(\|p(x_{i}^{T}QK^{T}x_{j})^{p-1}x_{i}x_{j}^{T}K\|^{2}_{F})$ (A.44)
-$=\sum_{i=1}^{N}\sum_{j=1}^{N}\mathop{\mathbb{E}}(p^{2}(x_{i}^{T}QK^{T}x_{j})^{2p-2}\sum_{k=1}^{D}\sum_{l=1}^{d}(x_{ik}\sum_{m=1}^{D}x_{jm}k_{ml})^{2}).$ (A.45)
+$\mathop{\mathbb{E}}(\|\frac{\partial(XQK^{T}X^{T})^{p}}{\partial Q}\|^{2}_{F})$ $=\sum_{i=1}^{N}\sum_{j=1}^{N}\mathop{\mathbb{E}}(\|\frac{(\partial x_{i}^{T}QK^{T}x_{j})^{p}}{\partial Q}\|^{2}_{F})$ (57)
+$=\sum_{i=1}^{N}\sum_{j=1}^{N}\mathop{\mathbb{E}}(\|p(x_{i}^{T}QK^{T}x_{j})^{p-1}\frac{\partial x_{i}^{T}QK^{T}x_{j}}{\partial Q}\|^{2}_{F})$ (58)
+$=\sum_{i=1}^{N}\sum_{j=1}^{N}\mathop{\mathbb{E}}(\|p(x_{i}^{T}QK^{T}x_{j})^{p-1}x_{i}x_{j}^{T}K\|^{2}_{F})$ (59)
+$=\sum_{i=1}^{N}\sum_{j=1}^{N}\mathop{\mathbb{E}}(p^{2}(x_{i}^{T}QK^{T}x_{j})^{2p-2}\sum_{k=1}^{D}\sum_{l=1}^{d}(x_{ik}\sum_{m=1}^{D}x_{jm}k_{ml})^{2}).$ (60)
 We know that
-$(x_{i}^{T}QK^{T}x_{j})^{2p-2}$ $=(\sum_{l=1}^{d}((\sum_{k=1}^{D}x_{ik}q_{kl})\cdot(\sum_{m=1}^{D}x_{jm}k_{ml})))^{2p-2}$ (A.46)
-$=(\sum_{l=1}^{d}\sum_{k=1}^{D}\sum_{m=1}^{D}x_{ik}q_{kl}x_{jm}k_{ml})^{2p-2}$ (A.47)
-$=(\sum_{k=1}^{D}\sum_{m=1}^{D}x_{ik}x_{jm}\sum_{l=1}^{d}q_{kl}k_{ml})^{2p-2}$ (A.48)
-$=(\sum_{k=1}^{D}\sum_{m=1}^{D}x_{ik}x_{jm}a_{km})^{2p-2},$ (A.49)
+$(x_{i}^{T}QK^{T}x_{j})^{2p-2}$ $=(\sum_{l=1}^{d}((\sum_{k=1}^{D}x_{ik}q_{kl})\cdot(\sum_{m=1}^{D}x_{jm}k_{ml})))^{2p-2}$ (61)
+$=(\sum_{l=1}^{d}\sum_{k=1}^{D}\sum_{m=1}^{D}x_{ik}q_{kl}x_{jm}k_{ml})^{2p-2}$ (62)
+$=(\sum_{k=1}^{D}\sum_{m=1}^{D}x_{ik}x_{jm}\sum_{l=1}^{d}q_{kl}k_{ml})^{2p-2}$ (63)
+$=(\sum_{k=1}^{D}\sum_{m=1}^{D}x_{ik}x_{jm}a_{km})^{2p-2},$ (64)
 where $a_{km}=\sum_{l=1}^{d}q_{kl}k_{ml}$. Let $z_{ij}=\sum_{k=1}^{D}\sum_{m=1}^{D}x_{ik}x_{jm}a_{km}$ Thus we have
-$\mathop{\mathbb{E}}(\|\frac{\partial(XQK^{T}X^{T})^{p}}{\partial Q}\|^{2}_{F})$ $=$ $p^{2}\sum_{i=1}^{N}\sum_{j=1}^{N}\mathop{\mathbb{E}}(z_{ij}^{2p-2}\sum_{k=1}^{D}\sum_{l=1}^{d}(x_{ik}\sum_{m=1}^{D}x_{jm}k_{ml})^{2})$ (A.50)
-$=$ $\sum_{i=1}^{N}\sum_{j=1}^{N}\sum_{k=1}^{D}\sum_{l=1}^{d}(\sum_{m=1}^{D}\mathop{\mathbb{E}}(z_{ij}^{2p-2}x_{ik}^{2}x_{jm}^{2}k_{ml}^{2})$ (A.51)
-$+\sum_{m=1}^{D}\sum_{n=1,n\neq m}^{D}\mathop{\mathbb{E}}(z_{ij}^{2p-2}x_{ik}^{2}x_{jm}k_{ml}x_{jn}k_{nl}))$ (A.52)
-$=$ $\sum_{i=1}^{N}\sum_{j=1}^{N}\sum_{k=1}^{D}\sum_{l=1}^{d}(\sum_{m=1}^{D}\mathop{\mathbb{E}}(z_{ij}^{2p-2}x_{ik}^{2}x_{jm}^{2}k_{ml}^{2})$ (A.53)
-$+\sum_{m=1}^{D}\sum_{n=1,n\neq m}^{D}\mathop{\mathbb{E}}(z_{ij}^{2p-3}x_{ik}^{2}x_{jm}^{2}k_{ml}^{2}x_{jn}^{2}k_{nl}^{2}))$ (A.54)
-$\approx$ $N^{2}Dd(D(D^{2p-2}d^{p-1}(2p-3)!!\sigma_{x}^{4p}\sigma_{w}^{4p-2})+0$ (A.55)
-$=$ $N^{2}D^{2p}d^{p}(2p-3)!!\sigma_{x}^{4p}\sigma_{w}^{4p-2}$ (A.56)
+$\mathop{\mathbb{E}}(\|\frac{\partial(XQK^{T}X^{T})^{p}}{\partial Q}\|^{2}_{F})$ $=p^{2}\sum_{i=1}^{N}\sum_{j=1}^{N}\mathop{\mathbb{E}}(z_{ij}^{2p-2}\sum_{k=1}^{D}\sum_{l=1}^{d}(x_{ik}\sum_{m=1}^{D}x_{jm}k_{ml})^{2})$ (65)
+$=\sum_{i=1}^{N}\sum_{j=1}^{N}\sum_{k=1}^{D}\sum_{l=1}^{d}(\sum_{m=1}^{D}\mathop{\mathbb{E}}(z_{ij}^{2p-2}x_{ik}^{2}x_{jm}^{2}k_{ml}^{2})$ (66)
+$+\sum_{m=1}^{D}\sum_{n=1,n\neq m}^{D}\mathop{\mathbb{E}}(z_{ij}^{2p-2}x_{ik}^{2}x_{jm}k_{ml}x_{jn}k_{nl}))$ (67)
+$=\sum_{i=1}^{N}\sum_{j=1}^{N}\sum_{k=1}^{D}\sum_{l=1}^{d}(\sum_{m=1}^{D}\mathop{\mathbb{E}}(z_{ij}^{2p-2}x_{ik}^{2}x_{jm}^{2}k_{ml}^{2})$ (68)
+$+\sum_{m=1}^{D}\sum_{n=1,n\neq m}^{D}\mathop{\mathbb{E}}(z_{ij}^{2p-3}x_{ik}^{2}x_{jm}^{2}k_{ml}^{2}x_{jn}^{2}k_{nl}^{2}))$ (69)
+$\approx N^{2}Dd(D(D^{2p-2}d^{p-1}(2p-3)!!\sigma_{x}^{4p}\sigma_{w}^{4p-2})+0$ (70)
+$=N^{2}D^{2p}d^{p}(2p-3)!!\sigma_{x}^{4p}\sigma_{w}^{4p-2}$ (71)
 showing that we can bound the gradient by a quantity of the form $\mathcal{O}(N)$ and the proof is complete.
 A.2 Experiments
-A.2.1 Ablations
-In this section we carry out ablations on the experiments we did in the main paper.
-Scale ablations:
-The theory we developed in section 4.2 suggested that we needed to scale our polynomial activations by $\frac{1}{\sqrt{N}}$ to obtain a complexity bound of $\mathcal{O}(\sqrt{N})$.
-In general, we could also scale the polynomial activations by $\mathcal{O}(\frac{1}{\sqrt{N}})$ to see if we can get a better accuracr.
-Figure 10 carries out an ablation on both the ViT-Base architecture and the ViT-Small architecture to see how different scales affect the accuracy for the $x^{3}$ activation.
-We found that in general a scale from $\frac{1}{8}$ to $\frac{1}{25}$ seemed to perform very well.
-Figure 10: We show how the top-1% accuracy changes as the scale of the activation $x^{3}$ for a ViT-Base (left) and ViT-Small (right) architecture.
-The x-axis plots the denominator of the scale for easier readability.
-In other words as we go to the right of the x-axis the scale used on the activation $x^{3}$ is getting smaller.
-Activation ablations:
-Our theory primarily compared the activations $\frac{1}{\sqrt{N}}x^{3}$ and $\frac{1}{\sqrt{N}}x$ with softmax where $N$ was the sequence length of the input.
-In this section we carry out further experimental comparisons of our activations with other activations.
-We compare with a variety of activations used in the literature as well as the exponential function $Exp=e^{-x}$ which can be thought of as softmax without the row normalization scaling.
-Table 5 shows the results of the experiments.
-Activation ViT-Base ViT-small
-softmax 79.6 80.2
-$\frac{x^{3}}{14}$ 79.6 80.5
-$\frac{x}{14}$ 76.9 77.8
-$\frac{x^{2}}{14}$ 78.7 79.9
-$ReLU$ 77.3 77.5
-$ELU$ 78.2 78.5
-$GELU$ 78.2 78.4
-$Tanh$ 77.1 77.3
-$Exp$ 77.9 78.0
-Table 5: Top-1% accuracy with various activations on a ViT-Base and ViT-Small architecture.
-A.2.2 Frobenius norm computations
-In section 5.1 we showed plots of the Frobenius norm of the self-attention matrix and for the Jacobian of the self-attention matrix for different scalings of the $x^{3}$ and $x$ activations along with softmax.
+A.2.1 Hardware
+The vision transformer experiments in Section 6.1, the object detection and instance segmentation experiments in Section 6.2 and the Nyströmformer experiments from Section 6.3 were all carried out on Nvidia A100 GPUs.
+The physics informed transformer experiments in Section A.2.5 were carried out on a Nvidia A6000 GPU.
+A.2.2 Experimental hyperparameters
+Vision transformers in Section 6.1.
+In Section 6.1 we tested four different vision transformers, ViT-B [10], DeiT-B [25], Swin-B [17] and XCiT-M [3], with the activations softmax, $x^{3}$ + fixed scaling, $x^{3}$ + dynamic scaling and $x^{3}$.
+The training strategy follow the exact strategy used in each of the original papers, we used the Timm [28] libraries to train our models.
+Object detection and instance segmentation in Section 6.2.
+Each of the models are trained for 36 epochs using the AdamW optimizer with learning rate of $10^{-4}$, $0.05$ weight decay and batch size of $16$.
+Nyströmformer in Section 6.3.
+For the Nyströmformer experiments we used the exact same training strategy form the original paper [30]. Model were trained using the GitHub provided by the authors of [30].
+A.2.3 Frobenius norm computations
+In Section 5 we showed plots of the Frobenius norm of the self-attention matrix and for the Jacobian of the self-attention matrix for softmax, $\frac{1}{14}x^{3}$, and $x^{3}$.
 This was done for a ViT-Tiny architecture on the Tiny-ImageNet dataset.
-Figure 11 shows the plots of the Frobenius norm of the self-attention matrix for the Tiny-ViT architecture, during training, for all layers averaged over the heads within each layer.
+Figure 6 shows the plots of the Frobenius norm of the self-attention matrix for the ViT-Tiny architecture, during training, for all layers averaged over the heads within each layer.
 Figure 7 shows the Frobenius norm of the Jacobian of the self-attention matrix during training for each layer, averaged over the total number of heads within each layer.
-Figures 13 and 14 show similar results for the ViT-Small architecture.
-Figure 11: Frobenius norm of self-attention matrix for different scaled activations on ViT-Tiny during training (zoom in for better viewing).
-Figure 12: Frobenius norm of Jacobian of self-attention matrix for different scaled activations on ViT-Tiny during training (zoom in for better viewing).
-Figure 13: Frobenius norm of self-attention matrix for different scaled activations on ViT-Small during training (zoom in for better viewing).
-Figure 14: Frobenius norm of Jacobian of self-attention matrix for different scaled activations on ViT-Small during training (zoom in for better viewing).
+Figure 6: Frobenius norm of self-attention matrix for softmax, $\frac{1}{16}x^{3}$ and $x^{3}$ on ViT-Tiny during training on the Tiny-ImageNet dataset (zoom in for better viewing).
+Figure 7: Frobenius norm of the Jacobian of the self-attention matrix for softmax, $\frac{1}{16}x^{3}$ and $x^{3}$ on ViT-Tiny during training on the Tiny-ImageNet dataset (zoom in for better viewing).
+A.2.4 Discussion on Linear attention and other polynomials on ViTs
+In the experiments Section 6 we showed our insights using the polynomial $x^{3}$ as this was a non-linear polynomial that did not satisfy the key properties that softmax did.
+In this section we will compare softmax with other polynomials that we also found to perform well.
+The results for ViTs is shown in Table 5.
+As we can see from that table each polynomial does far better when scaled using the theory from Section 4.
+We note that when we tried to train polynomial higher than order $6$ they did not train well.
+On inspecting the weights we found that several were very close to zero leading to a gradient vanishing problem.
+We hypothesize that this is because the transformers were randomly initialized with weights about zero.
+Thus taking such a weight and applying a polynomial of the form $x^{k}$ with $k\geq 6$ would make those weights orders of magnitude smaller, making it difficult to train after some point.
+Table 5: Comparisons of pre-training models with different activation functions on ImageNet-1k. We report top-1 classification accuracy (%).
+Models
+ViT-B DeiT-B Swin-B XciT-M
+softmax 80.3 81.5 83.5 82.7
+$x$ + fixed scale 78.4 79.4 80.6 80.2
+$x$ + dynamic scale 78.7 79.5 80.7 80.4
+$x$ 74.1 77.9 78.3 78.1
+$x^{2}$ + fixed scale 80.1 81.5 83.4 82.5
+$x^{2}$ + dynamic scale 80.3 81.6 83.5 82.7
+$x^{2}$ 77.8 78.2 79.8 79.4
+$x^{4}$ + fixed scale 80.3 81.5 83.7 82.7
+$x^{4}$ + dynamic scale 80.3 81.6 83.7 82.8
+$x^{4}$ 77.9 78.6 79.9 79.6
+$x^{5}$ + fixed scale 80.3 81.4 83.4 82.5
+$x^{5}$ + dynamic scale 80.3 81.5 83.4 82.6
+$x^{5}$ 77.7 78.0 79.4 79.5
+A.2.5 Physics informed transformers
+In [31] transformers for modeling partial differential equations (PDEs) were introduced and shown to yield superior performance on a variety of PDEs.
+In this section we compare softmax with the polynomial activation we have been using in the previous section on two common PDEs within the literature, namely the convection PDE and the 1d-Reaction PDE.
+Convection PDE:
+This is a 1-dimensional PDE that models transport phenomena. The PDE system is defined by the following equations
+$\frac{\partial u}{\partial t}+\beta\frac{\partial u}{\partial x}=0,\text{ for }x\in[0,2\pi],t\in[0,1]$ (72)
+$IC:u(x,0)=sin(x),BC:u(0,t)=u(2\pi,t)$ (73)
+where $\beta$ is the convection coefficient and is set as $\beta=50$. IC stands for initial condition and BC for boundary condition. For more details on this system consult [31].
+1d-reaction PDE:
+The 1d-reaction equation is commonly used in modeling chemical reactions. It is defined by the following system.
+$\frac{\partial u}{\partial t}-\rho u(1-u)=0\text{ for }x\in[0,2\pi],t\in[0,1]$ (74)
+$IC:u(x,0)=exp\big{(}\frac{(x-\pi)^{2}}{2(\pi/4)^{2}}\big{)},BC:u(0,t)=u(2\pi,t)$ (75)
+where $\rho$ is known as the reaction coefficient and is set as $\rho=5$. For more details on this system consult [31].
+As in [31] we considered a physics informed transformer (PINNsformer) on the above two PDEs.
+We compared softmax with $x^{3}$ + fixed scaling, $x^{3}$ + dynamic scaling, where the scale was initialized to $\frac{1}{\sqrt{51}}$ as the input sequence length is $51$, and $x^{3}$.
+As in [31] we computed the training loss, and for testing the relative mean absolute error (rMAE), and the relative mean squared error (rMSE).
+Table 6 shows the results from which it can be seen that both $x^{3}$ + fixed scaling and $x^{3}$ + dynamic scaling are competitive with the original softmax.
+Table 6: Comparison of different activations on PINNsformer for Convection and 1D-Reaction problems.
+Convection PDE 1D-Reaction PDE
+Loss rMAE rMSE Loss rMAE rMSE
+softmax 5.4e-5 4.1e-2 4.4e-2 3.8e-6 3.1e-2 5.8e-2
+$x^{3}$ + fixed scale 4.9e-5 3.9e-2 4.3e-2 3.1e-6 3.0e-2 5.6e-2
+$x^{3}$ + dynamic scale 4.1e-5 3.5e-2 4.0e-2 2.8e-6 2.7e-2 5.4e-2
+$x^{3}$ 7.2e-5 9.8e-2 1.1e-1 6.2e-6 6.1e-2 9.8e-2
+The training of the physics informed transformers followed the exact recipe in the original paper [31] with their data and training schemes available from their GitHub [1].
+The evaluation metrics are defined by
+$\text{rMAE}=\frac{\sum_{n=1}^{N}\left|\hat{u}(x_{n},t_{n})-u(x_{n},t_{n})\right|}{\sum_{n=1}^{N}\left|u(x_{n},t_{n})\right|}$ (76)
+$\text{rMSE}=\sqrt{\frac{\sum_{n=1}^{N}\left|\hat{u}(x_{n},t_{n})-u(x_{n},t_{n})\right|^{2}}{\sum_{n=1}^{N}\left|u(x_{n},t_{n})\right|^{2}}}$ (77)
+where $N$ is the number of testing points, $\hat{u}$ is the neural network approximation, and $u$ is the ground truth.

### v3

@@
 For the theoretical analysis in this paper we will only use the dot-product similarity $qk^{T}$ and call the $N\times N$ matrix $softmax(qk^{T})$ the softmax self-attention matrix.
-In the experiments, Section 6, we will empirically validate our theoretical framework on more general softmax attention blocks used in state of the art transformers such as DeiT [25], Swin Transformer [17] and XciT [30].
+In the experiments, Section˜6, we will empirically validate our theoretical framework on more general softmax attention blocks used in state of the art transformers such as DeiT [25], Swin Transformer [17] and XciT [30].
 For general transformer architectures, multiple heads $\mathbf{A}_{i}$ for $1\leq i\leq n$ are used.
@@
 where $\mathbf{Jsoftmax}(A)$ is the Jacobian of $\mathbf{softmax}$ at $A$.
-Given a matrix $A\in\mathbb{R}^{n\times m}$, we denote its Frobenius norm by $||A||_{F}$.
+Given a matrix $A\in\mathbb{R}^{n\times m}$, we denote its Frobenius norm by $A||_{F}$.
 Additionally, we use the notation $\mathbb{E}$ to represent the expectation of a random variable, where the specific random variable being considered will be clear from the context.
@@
 We then have the following bounds on the Frobenius norms
-$||\mathbf{softmax}(A)||_{F}$ $\leq\sqrt{N}$ (6)
-$||\mathbf{\nabla softmax}(A)||_{F}$ $\leq 2\sqrt{N}.$ (7)
+[…]
+[…]
 The key implication of theorem 4.1 is that during the training of a transformer with softmax self-attention, the Frobenius norm of each softmax self-attention matrix remains bounded by a value that grows as $\mathcal{O}(\sqrt{N})$.
@@
 The proof hinges on the fact that the row normalization inherent in softmax effectively controls the Frobenius norm.
-For a detailed proof see Section A.1.1.
+For a detailed proof see Section˜A.1.1.
 4.2 Polynomial activations for self-attention
-In Section 4.1, we demonstrated that softmax implicitly regularizes the Frobenius norm of the self-attention matrix.
+In Section˜4.1, we demonstrated that softmax implicitly regularizes the Frobenius norm of the self-attention matrix.
 Building on this, we now show that by scaling specific polynomial activations, a similar regularization effect on the Frobenius norm can be achieved in expectation, closely replicating the impact of softmax.
@@
 We have the following expectations of the Frobenius norms of powers of the $N\times N$ matrix $(XQK^{T}X^{T})^{p}$ for $p\geq 1$
-$\mathbb{E}\bigg{|}\bigg{|}\bigg{(}\frac{XQK^{T}X^{T}}{\sqrt{d}}\bigg{)}^{p}\bigg{|}\bigg{|}_{F}\leq\mathcal{O}(N)$ (8)
+[…]
 By scaling such an activation by $\frac{1}{\sqrt{N}}$ we can obtain a $\mathcal{O}(\sqrt{N})$ bound.
@@
 Assume the same conditions as in theorem 4.2. Then
-$\mathbb{E}\bigg{|}\bigg{|}\frac{1}{\sqrt{N}}\bigg{(}\frac{XQK^{T}X^{T}}{\sqrt{d}}\bigg{)}^{p}\bigg{|}\bigg{|}_{F}\leq\mathcal{O}(\sqrt{N}).$ (9)
-Corollary 4.3 establishes that activations of the form $\phi(x):=\frac{1}{\sqrt{N}}x^{p}$ provide a level of regularization, in expectation, similar to that of softmax when applied to the self-attention matrix.
+[…]
+Corollary˜4.3 establishes that activations of the form $\phi(x):=\frac{1}{\sqrt{N}}x^{p}$ provide a level of regularization, in expectation, similar to that of softmax when applied to the self-attention matrix.
 The proof of theorem 4.2 can be found in appendix A.1.2.
-The next property we want to prove is one similar to the gradient bound obtained in Theorem 4.1.
+The next property we want to prove is one similar to the gradient bound obtained in Theorem˜4.1.
 Since the self-attention matrix has parameters given by the queries $Q$ and keys $K$ [26], this implies that during the training of a transformer the $Q$ and $K$ matrices are the only aspects of the self-attention matrix that get updated.
@@
 Then the expectation of the of the derivative of the matrix $\frac{(XQK^{T}X^{T})^{p}}{\sqrt{d}}$ w.r.t the $Q$ parameter matrix for $p\geq 1$ is given by
-$\mathbb{E}\bigg{|}\bigg{|}\frac{\partial}{\partial Q}\bigg{(}\frac{(XQK^{T}X^{T})^{p}}{\sqrt{d}}\bigg{)}\bigg{|}\bigg{|}\leq\mathcal{O}(N)$ (10)
+[…]
 Figure 1: Training ViT-Tiny with the activation $\phi(x)=x^{3}$ with different sequence lengths and different scales.
@@
 Assume the same condition as in theorem 4.4. Then
-$\mathbb{E}\bigg{|}\bigg{|}\frac{1}{\sqrt{N}}\frac{\partial}{\partial Q}\bigg{(}\frac{(XQK^{T}X^{T})^{p}}{\sqrt{d}}\bigg{)}\bigg{|}\bigg{|}\leq\mathcal{O}(\sqrt{N}).$ (11)
+[…]
 An analogous estimate holds for derivatives with respect to the $K$ matrix. The proof of theorem 4.4 can be found in appendix A.1.2.
 Remark 4.6.
-Corollary 4.3 and Corollary 4.5 suggest that polynomial activations of the form $\phi(x)=\frac{1}{\sqrt{N}}x^{p}$, with $p>0$, can achieve performance comparable to softmax when applied to self-attention matrices.
+Corollary˜4.3 and Corollary˜4.5 suggest that polynomial activations of the form $\phi(x)=\frac{1}{\sqrt{N}}x^{p}$, with $p>0$, can achieve performance comparable to softmax when applied to self-attention matrices.
 We point out that both corollaries rested on the assumption that $X$, $Q$ and $K$ are i.i.d random variables.
 In general, this is only true at initialization when training a transformer [2].
-Although this is a limitation in the theory the experiments in Section 6 show that this insight can be used effectively to develop new attention blocks that perform comparable to softmax yet violate the three conditions of positivity, normalized rows summing to $1$ and sparsity showing that attention blocks do not need to be modeled as a probability distribution.
+Although this is a limitation in the theory the experiments in Section˜6 show that this insight can be used effectively to develop new attention blocks that perform comparable to softmax yet violate the three conditions of positivity, normalized rows summing to $1$ and sparsity showing that attention blocks do not need to be modeled as a probability distribution.
 5 Testing the theory
-In this section we test the theory developed in Section 4.2 on small vision transformers.
+In this section we test the theory developed in Section˜4.2 on small vision transformers.
 We will consider the activation $\phi(x)=\frac{1}{k}x^{3}$, where $k>0$ is a fixed scale, as this activation clearly violates the three key conditions of softmax based attention; positivity, normalization and sparsity.
-We found that polynomials $\frac{1}{k}x^{p}$ for $p>3$ did not perform well during training as they witnessed a gradient vanishing problem due to the fact that the function $x^{p}$ for $p$ large have very small values around 00.
+We found that polynomials $\frac{1}{k}x^{p}$ for $p>3$ did not perform well during training as they witnessed a gradient vanishing problem due to the fact that the function $x^{p}$ for $p$ large have very small values around 0.
 The first experiment we conducted was to test how the Top-$1\%$ accuracy changes for a ViT-Tiny vision transformer [22], trained on Tiny-Imagenet dataset [21], as we change the sequence length $N$ of the input and the scale predicted in corollaries 4.3 and 4.5 when using the activation $\frac{1}{k}x^{3}$.
@@
 Thus the best accuracy should occur when $k=\mathcal{O}(\sqrt{N})$ and should degrade for other values due to training instability.
-Figure 1 shows the results of the experiment, we note that the $x-axis$ plots the values of $k$.
+Figure˜1 shows the results of the experiment, we note that the xaxisx-axis plots the values of $k$.
 We see that as the sequence length increases the factor of $k$ needs to increase so that the activation $\frac{1}{k}x^{3}$ performs well on the ViT-Tiny architecture as predicted by the theory developed in corollaries 4.3 and 4.5.
 In a second experiment we decided to compare the performance of the original ViT-Tiny architecture, that uses an input sequence length of $256$, with a softmax activation and a polynomial activation on the Tiny-ImageNet dataset.
-In Corollary 4.3 and Corollary 4.5 it was pointed out that the scaling of the polynomial is important to keep the Frobenius norm of the polynomial attention matrix from becoming too large.
+In Corollary˜4.3 and Corollary˜4.5 it was pointed out that the scaling of the polynomial is important to keep the Frobenius norm of the polynomial attention matrix from becoming too large.
 The results of those corollaries suggested that a scale to use is $\frac{1}{\sqrt{N}}$, $N$ being the sequence length, which in this case is $\frac{1}{\sqrt{256}}=\frac{1}{16}$.
@@
 To begin with we computed the Frobenius norm of the attention matrix throughout training averaged over all the heads in layers 2, 7 and 12 of the ViT-Tiny architecture when trained on the Tiny-ImageNet dataset.
-Figure 2 plots the results.
+Figure˜2 plots the results.
 We see from the figure that the Frobenius norm of of the $x^{3}$ archicture is much larger than softmax but scaling it by $\frac{1}{16}$ brings it down to softmax levels.
-Similarly, Figure 3 shows the Jacobian’s Frobenius norm, where scaling also brings the norms closer to softmax, ensuring more stable gradients.
-Further plots for other layers are in Section A.2.3.
-Table 1 presents the final Top-1% accuracy achieved by each activation function.
+Similarly, Figure˜3 shows the Jacobian’s Frobenius norm, where scaling also brings the norms closer to softmax, ensuring more stable gradients.
+Further plots for other layers are in Section˜A.2.3.
+Table˜1 presents the final Top-1% accuracy achieved by each activation function.
 Notably, $\frac{x^{3}}{16}$ delivers the best performance.
@@
 The goal is to empirically challenge the conventional softmax properties, positivity, row-normalization, and sparsity—and examine whether these conditions are truly necessary for good transformer performance.
-Building on the theoretical foundations from Section 4, we focus on the cubic polynomial $x^{3}$ as a test case.
+Building on the theoretical foundations from Section˜4, we focus on the cubic polynomial $x^{3}$ as a test case.
 Notably, $x^{3}$ introduces both positive and negative values, does not normalize rows to sum to 1, and generally produces dense attention matrices—violating all three traditional softmax conditions.
@@
   • 1.
-    Fixed scale: Following Section 4, we scale $x^{3}$ by the inverse square root of the sequence length (which is fixed throughout training), as theory suggests this maintains optimization stability.
+    Fixed scale: Following Section˜4, we scale $x^{3}$ by the inverse square root of the sequence length (which is fixed throughout training), as theory suggests this maintains optimization stability.
   • 2.
     Dynamic scale: Recognizing that the assumption of i.i.d.
-normal-distributed $Q$, $K$, $V$, used in Section 4, holds primarily at initialization, we also explore a learnable scale.
+normal-distributed $Q$, $K$, $V$, used in Section˜4, holds primarily at initialization, we also explore a learnable scale.
 This dynamic scale is initialized as above but optimized during training.
@@
 Each transformer was trained following the approach in each paper.
-The results are shown in Table 2.
+The results are shown in Table˜2.
 As can be seen from the table the $x^{3}$ activation with a dynamic scale performed the best and the one with a fixed scale performed comparable to softmax.
-On the other hand $x^{3}$ (with no scale) underperformed on all ViTs showing how important the theory on scaling as developed in Section 4 is.
-For a discussion on other polynomials and linear attention see Section A.2.4.
+On the other hand $x^{3}$ (with no scale) underperformed on all ViTs showing how important the theory on scaling as developed in Section˜4 is.
+For a discussion on other polynomials and linear attention see Section˜A.2.4.
 Table 2: Comparisons of pre-training models with different activation functions on ImageNet-1k. We report top-1 classification accuracy (%).
@@
 We visualized heatmaps of the attention matrices after convergence, focusing on two representative layers and heads, and averaged over a fixed batch of 128 samples.
-Figure 4 shows results for layer 2, head 8, where the $x^{3}$ + dynamic scale activation produces attention scores with both positive and negative values, unlike softmax.
-Similarly, Figure 5 illustrates distinct patterns for layer 12, head 6, highlighting how the two activations differ in learned attention distributions.
+Figure˜4 shows results for layer 2, head 8, where the $x^{3}$ + dynamic scale activation produces attention scores with both positive and negative values, unlike softmax.
+Similarly, Figure˜5 illustrates distinct patterns for layer 12, head 6, highlighting how the two activations differ in learned attention distributions.
 Interpretability.
 Even with these differences, the $x^{3}$ + dynamic scale activation still learns meaningful attention patterns.
-For instance, in Figure 4 (left), we observe 14 bands about the diagonal, indicating that the head has learned to attend to patches in the same image row as the query patch, sufficient for effective image classification.
+For instance, in Figure˜4 (left), we observe 14 bands about the diagonal, indicating that the head has learned to attend to patches in the same image row as the query patch, sufficient for effective image classification.
 This suggests that the sign of attention values (positive or negative) is not inherently critical for attention allocation.
-Further, Figure 5 (left) reveals vertical lines, showing attention that depends solely on key positions, independent of the query position.
+Further, Figure˜5 (left) reveals vertical lines, showing attention that depends solely on key positions, independent of the query position.
 Smaller key indices receive higher attention weights, focusing model capacity where it matters most for classification tasks.
@@
 We conducted experiments on XCiT-S12 models using 16×16 patches with the activations $x^{3}$ + fixed scale, $x^{3}$ + dynamic scale, with scale initialized to $\frac{1}{14}$, and softmax.
-We found when we trained with $x^{3}$ the model would not converge thereby suggesting that a scaling is extremely important as predicted by the theory in Section 4.2.
-The training hyperparameters are given in Section A.2.2.
-Table 3 shows the results of the experiment.
+We found when we trained with $x^{3}$ the model would not converge thereby suggesting that a scaling is extremely important as predicted by the theory in Section˜4.2.
+The training hyperparameters are given in Section˜A.2.2.
+Table˜3 shows the results of the experiment.
 From the table we see that the activation $x^{3}$ + dynamic scale obtains better performance to softmax with $x^{3}$ + fixed scale obtaining comparable performance.
@@
 We followed the training regime of the original paper [30].
-The results are shown in Table 4.
+The results are shown in Table˜4.
 As can be seen from the table the $x^{3}$ + dynamic scale activation results in better performance compared to softmax with $x^{3}$ + fixed scale performing comparably.
@@
 These models effectively capture interactions between spatial and temporal components inherent in dynamical systems.
-In Section A.2.5, we extend our polynomial attention framework to Physics-Informed Transformers, showing that polynomial activations can outperform softmax in this setting.
+In Section˜A.2.5, we extend our polynomial attention framework to Physics-Informed Transformers, showing that polynomial activations can outperform softmax in this setting.
 8 Limitations
@@
 Through extensive experiments across a variety of transformer applications, we demonstrated that these alternative activations not only compete with but can outperform softmax, offering a new perspective on attention mechanisms in transformers.
+10 Acknowledgements
+Simon Lucey acknowledges support from the Australian Research Council (ARC) through the Discovery Project DP220103803.
 Appendix A Appendix
@@
 By definition of the Frobenius norm we then see that
-$||\mathbf{softmax}(A)||_{F}^{2}$ $=\bigg{(}\frac{1}{\sum_{j=1}^{n}e^{a_{1j}}}\bigg{)}^{2}(e^{2a_{11}}+\cdots e^{2a_{11}})+\cdots$ (13)
+[…]
 $\hskip 28.45274pt+\bigg(\frac{1}{\sum_{j=1}^{n}e^{a_{Nj}}}\bigg)^{2}(e^{2a_{N1}}+\cdots e^{2a_{NN}})$ (14)
@@
 where the second inequality uses the fact that for non-negative numbers $a$ and $b$ we always have that $a^{2}+b^{2}\leq(a+b)^{2}$.
-It then immediately follows that $||\mathbf{softmax}(A)||_{F}\leq\sqrt{N}$ and this proves the first inequality in the statement of theorem 4.1.
+It then immediately follows that $\mathbf{softmax}(A)||_{F}\leq\sqrt{N}$ and this proves the first inequality in the statement of theorem 4.1.
 We move on to prove the second inequality in the statement of theorem 4.1. For this, let us write each entry of the matrix on the right of (12) as follows:
@@
 From these computations we see that
-$||\nabla\mathbf{softmax}(A)||_{F}^{2}=||\nabla F_{11}||_{F}^{2}+\cdots+||\nabla F_{NN}||_{F}^{2}.$ (25)
-We will proceed by bounding each collection $||\nabla F_{i1}||_{F}^{2}+\cdots+||\nabla F_{1N}||_{F}^{2}$ separately then add up all the bounds. We have
-$||\nabla F_{i1}||_{F}^{2}+\cdots+||\nabla F_{1N}||_{F}^{2}$ $=|F_{i1}-F_{i1}^{2}|^{2}+|F_{i1}F_{i2}|^{2}+\cdots+|F_{i1}F_{iN}|^{2}$ (26)
+[…]
+[…]
+[…]
 $\hskip 14.22636pt+|F_{i2}F_{i1}|^{2}+|F_{i2}-F_{i2}^{2}|^{2}+\cdots+|F_{i2}F_{iN}|^{2}$ (27)
-[…]
+$\hskip 14.22636pt+\cdots\cdots\cdots\cdots+$ (28)
 $\hskip 14.22636pt+|F_{iN}F_{i1}|^{2}+|F_{iN}F_{i2}|^{2}+\cdots+|F_{iN}-F_{iN}^{2}|^{2}$ (29)
@@
 $\hskip 14.22636pt(F_{i2})^{2}(|F_{i1}|+|1-F_{i2}|+\cdots+|F_{iN}|)^{2}$ (31)
-[…]
+$\hskip 14.22636pt+\cdots\cdots\cdots\cdots+$ (32)
 $\hskip 14.22636pt+(F_{iN})^{2}(|F_{i1}|+|F_{i2}|+\cdots+|1-F_{iN}|)^{2}.$ (33)
@@
 This means we get the bound
-$||\nabla F_{i1}||_{F}^{2}+\cdots+||\nabla F_{1N}||_{F}^{2}$ $\leq 4F_{i1}^{2}(\widehat{F_{i1}}+F_{i2}+\cdots+F_{iN})$ (34)
 […]
-$\hskip 14.22636pt+4F_{iN}^{2}(F_{i1}+F_{i2}+\cdots+\widehat{F_{iN}})$ (36)
+$\hskip 14.22636pt+\cdots\cdots\cdots\cdots+$ (35)
+[…]
 $\leq 4(F_{i1}^{2}+\cdots F_{iN}^{2})$ (37)
 $=4.$ (38)
-Putting all the bounds together for each of the terms $N$ terms $||\nabla F_{i1}||_{F}^{2}+\cdots+||\nabla F_{1N}||_{F}^{2}$ we get
-$||\nabla\mathbf{softmax}(A)||_{F}^{2}\leq 4N$ (39)
+[…]
+$\nabla\mathbf{softmax}(A)||_{F}^{2}\leq 4N$ (39)
 and this implies
-$||\nabla\mathbf{softmax}(A)||_{F}\leq 2\sqrt{N}.$ (40)
+$\nabla\mathbf{softmax}(A)||_{F}\leq 2\sqrt{N}.$ (40)
 This finishes the proof of theorem 4.1.
 A.1.2 Proofs for theorems section 4.2
-In this section we will give the proof of theorems 4.2 and 4.4.
+In this section we will give the proof of theorems 4.2 and $4.$
 Proof of theorem 4.2.
@@
 For the case that $p>1$ we proceed in a similar way.
-The key observation is that odd powers, in the matrix expansion, will have expectaion 00, so we need only consider the even powers.
+The key observation is that odd powers, in the matrix expansion, will have expectaion 0, so we need only consider the even powers.
 Therefore, suppose $\mathbf{C}=(\mathbf{A}\mathbf{B})^{p}$.
@@
 $=\mathop{\mathbb{E}}(\sum_{i=1}^{D}a_{1i}^{2p}b_{i1}^{2p}+\sum_{i=1}^{D}\sum_{j=1,j\neq i}^{D}a_{1i}^{2p-2}b_{i1}^{2p-2}a_{1j}^{2}b_{j1}^{2}+\cdots).$
-Note that the first term only has a count of $D$ and the second term has a count of $D(D-1)$.
+Note that the first term only has a count of $D$ and the second term has a count of D(D1)D(D-1).
 Thus, we only need to consider the $\mathcal{O}(D^{p})$ term where all the components have a power of 2.
@@
 $=\frac{D!}{(D-p)!}\frac{2p!}{2p!!}\sigma_{1}^{2p}\sigma_{2}^{2p}$
-$=\frac{D!}{(D-p)!}(2p-1)!!\sigma_{1}^{2p}\sigma_{2}^{2p}.$
+$=\frac{D!}{(D-p)!}(2p-1)!\sigma_{1}^{2p}\sigma_{2}^{2p}.$
 $\frac{D!}{(D-p)!}$ can always be bounded above by $D^{p}$, so the expectation can be upper bounded by $D^{p}(2p-1)!!\sigma_{1}^{2p}\sigma_{2}^{2p}$ and thus we get a quantity of the form $\mathcal{O}(N)$.
@@
 A.2.1 Hardware
-The vision transformer experiments in Section 6.1, the object detection and instance segmentation experiments in Section 6.2 and the Nyströmformer experiments from Section 6.3 were all carried out on Nvidia A100 GPUs.
-The physics informed transformer experiments in Section A.2.5 were carried out on a Nvidia A6000 GPU.
+The vision transformer experiments in Section˜6.1, the object detection and instance segmentation experiments in Section˜6.2 and the Nyströmformer experiments from Section˜6.3 were all carried out on Nvidia A100 GPUs.
+The physics informed transformer experiments in Section˜A.2.5 were carried out on a Nvidia A6000 GPU.
 A.2.2 Experimental hyperparameters
-Vision transformers in Section 6.1.
-In Section 6.1 we tested four different vision transformers, ViT-B [10], DeiT-B [25], Swin-B [17] and XCiT-M [3], with the activations softmax, $x^{3}$ + fixed scaling, $x^{3}$ + dynamic scaling and $x^{3}$.
+Vision transformers in Section˜6.1.
+In Section˜6.1 we tested four different vision transformers, ViT-B [10], DeiT-B [25], Swin-B [17] and XCiT-M [3], with the activations softmax, $x^{3}$ + fixed scaling, $x^{3}$ + dynamic scaling and $x^{3}$.
 The training strategy follow the exact strategy used in each of the original papers, we used the Timm [28] libraries to train our models.
-Object detection and instance segmentation in Section 6.2.
+Object detection and instance segmentation in Section˜6.2.
 Each of the models are trained for 36 epochs using the AdamW optimizer with learning rate of $10^{-4}$, $0.05$ weight decay and batch size of $16$.
-Nyströmformer in Section 6.3.
+Nyströmformer in Section˜6.3.
 For the Nyströmformer experiments we used the exact same training strategy form the original paper [30]. Model were trained using the GitHub provided by the authors of [30].
 A.2.3 Frobenius norm computations
-In Section 5 we showed plots of the Frobenius norm of the self-attention matrix and for the Jacobian of the self-attention matrix for softmax, $\frac{1}{14}x^{3}$, and $x^{3}$.
+In Section˜5 we showed plots of the Frobenius norm of the self-attention matrix and for the Jacobian of the self-attention matrix for softmax, $\frac{1}{14}x^{3}$, and $x^{3}$.
 This was done for a ViT-Tiny architecture on the Tiny-ImageNet dataset.
-Figure 6 shows the plots of the Frobenius norm of the self-attention matrix for the ViT-Tiny architecture, during training, for all layers averaged over the heads within each layer.
-Figure 7 shows the Frobenius norm of the Jacobian of the self-attention matrix during training for each layer, averaged over the total number of heads within each layer.
+Figure˜6 shows the plots of the Frobenius norm of the self-attention matrix for the ViT-Tiny architecture, during training, for all layers averaged over the heads within each layer.
+Figure˜7 shows the Frobenius norm of the Jacobian of the self-attention matrix during training for each layer, averaged over the total number of heads within each layer.
 Figure 6: Frobenius norm of self-attention matrix for softmax, $\frac{1}{16}x^{3}$ and $x^{3}$ on ViT-Tiny during training on the Tiny-ImageNet dataset (zoom in for better viewing).
@@
 A.2.4 Discussion on Linear attention and other polynomials on ViTs
-In the experiments Section 6 we showed our insights using the polynomial $x^{3}$ as this was a non-linear polynomial that did not satisfy the key properties that softmax did.
+In the experiments Section˜6 we showed our insights using the polynomial $x^{3}$ as this was a non-linear polynomial that did not satisfy the key properties that softmax did.
 In this section we will compare softmax with other polynomials that we also found to perform well.
-The results for ViTs is shown in Table 5.
-As we can see from that table each polynomial does far better when scaled using the theory from Section 4.
+The results for ViTs is shown in Table˜5.
+As we can see from that table each polynomial does far better when scaled using the theory from Section˜4.
 We note that when we tried to train polynomial higher than order $6$ they did not train well.
@@
 As in [31] we computed the training loss, and for testing the relative mean absolute error (rMAE), and the relative mean squared error (rMSE).
-Table 6 shows the results from which it can be seen that both $x^{3}$ + fixed scaling and $x^{3}$ + dynamic scaling are competitive with the original softmax.
+Table˜6 shows the results from which it can be seen that both $x^{3}$ + fixed scaling and $x^{3}$ + dynamic scaling are competitive with the original softmax.
 Table 6: Comparison of different activations on PINNsformer for Convection and 1D-Reaction problems.